### v1

# SADG: Segment Any Dynamic Gaussian Without Object Trackers

Abstract
Understanding dynamic 3D scenes is fundamental for various applications, including extended reality (XR) and autonomous driving. Effectively integrating semantic information into 3D reconstruction enables holistic representation that opens opportunities for immersive and interactive applications. We introduce SADG, Segment Any Dynamic Gaussian Without Object Trackers, a novel approach that combines dynamic Gaussian Splatting representation and semantic information without reliance on object IDs. In contrast to existing works, we do not rely on supervision based on object identities to enable consistent segmentation of dynamic 3D objects. To this end, we propose to learn semantically-aware features by leveraging masks generated from the Segment Anything Model (SAM) and utilizing our novel contrastive learning objective based on hard pixel mining. The learned Gaussian features can be effectively clustered without further post-processing. This enables fast computation for further object-level editing, such as object removal, composition, and style transfer by manipulating the Gaussians in the scene. We further extend several dynamic novel-view datasets with segmentation benchmarks to enable testing of learned feature fields from unseen viewpoints. We evaluate SADG on proposed benchmarks and demonstrate the superior performance of our approach in segmenting objects within dynamic scenes along with its effectiveness for further downstream editing tasks. Our project page is available at: https://yunjinli.github.io/project-sadg/.
1 Introduction
We live in a dynamic, three-dimensional world. Most of the time, we describe our environment at a high level of objects. As we move around and observe the scene from different viewpoints, we complete the geometry and perceive the colors contributing to our object representation. At the same time, we infer dynamic properties with object motion and deformations. Despite pose, color, and shape changes, the underlying semantic meaning stays constant. This intuition can serve as a strong constraint for neural 3D representations that have gained recent popularity and promote innovation in augmented reality, gaming, and autonomous driving interactive applications. In these technologies, seamlessly integrating semantic information into underlying representation ensures an immersive experience and gives users advanced control over the dynamic scene contents.
Neural Radiance Fields (NeRFs) have revolutionized the field of 3D reconstruction and demonstrated excellent performance in synthesizing photorealistic novel views of a scene from any angle. The limitations of vanilla NeRFs promoted many follow-up works, including those that encompass semantic information by lifting 2D segmentation masks [24, 7, 19] or distilling features [21, 18] and those that faithfully represent dynamic scenes [41, 35, 27]. Due to continuous implicit representation, these methods are computationally intensive as they require re-rendering from multiple views to ensure edit consistency, which hinders their applicability for interactive downstream tasks.
Building on the strengths of NeRFs, Gaussian splatting (3DGS) [17] has introduced a powerful new approach that prioritizes efficiency and speed. Dynamic versions of 3DGS [30, 56] have offered faithful reconstruction of scenes with motion. These models can serve as a geometric prior for learning semantic features that are consistent in both space and time and open new possibilities for real-time interactive applications like scene editing and dynamic interactions. Towards this, a line of works has extended geometric properties of colored and explicit Gaussians with semantic information [6, 57], producing consistent novel view synthesis and segmentation. Nonetheless, these works are limited to static scenes and cannot handle objects in motion. There are only a few concurrent works [16, 25], which unify semantics and dynamics to comprehensively understand real-world environments. However, SA4D [16] relies on the supervision of object trackers and suffers from semantic identity conflicts in multi-view scenes. Moreover, their method is sensitive to mask noise and requires post-processing to remove degenerate Gaussians. At the same time, DGD [25] has a higher memory requirement and longer training time as they are learning 4D feature fields based on 512-dimensional CLIP [42] or 384-dimensional DINOv2 [34]. Our work aims to tackle both single- and multi-view representations with fast rendering times, minimal post-processing, and low memory footprint.
To this end, we introduce our novel framework, Segment Any Dynamic Gaussian Without Object Trackers (SADG), which effectively combines dynamic Gaussian splatting backbone [56] and semantic information without any reliance on trackers that maintain consistent object identities. To learn the spatio-temporal semantic field, we define a new contrastive objective based on hard positive and negative mining with the Segment Anything Masks (SAM) [20]. We inherit the strengths of zero-shot 2D scene understanding of the foundation model to learn the semantically-aware feature field and produce consistent spatio-temporal segmentation. Our 32-dimensional features are compact and do not depend on features of a 2D segmentation model. Rendering time and storage requirements are almost unaffected, enabling real-time interaction and editing. The learned feature field can be effectively clustered without any further post-processing and utilized in a range of downstream editing tasks, including object style transfer [14], re-coloring, composition, and removal. Given an early stage of development and benchmarks for semantic segmentation for dynamic scenes, we propose to extend existing novel-view synthesis datasets with semantic benchmarks. The proposed evaluation protocol includes single- and multi-view scenes with challenging motions and diverse contents, allowing extensive testing of semantically-aware latent representations from unseen viewpoints.
Our contributions can be summarized as follows:
We propose Segment Any Dynamic Gaussian Without Object Trackers (SADG), a new approach that achieves multi-view consistent segmentation of dynamic scenes without any tracking supervision.
We leverage information from 2D masks to learn semantically-aware latent representation for dynamic scenes using a novel contrastive learning objective.
We extensively evaluate our method on single- and multi-view scenes across five dynamic novel-view benchmarks and demonstrate state-of-the-art segmentation performance.
We demonstrate the general nature of our feature space and apply it to several downstream tasks, including object removal, style transfer, and scene composition.
We offer an interactive framework that allows to edit the scene with a simple mouse click or a text prompt. Our tool does not depend on input IDs or 2D strokes and operates in real-time.
2 Related Works
2.1 Semantic Feature Learning
With the emergence of various vision foundation models, such as Segment Anything Model (SAM) [20], DINO [5], and DINOv2 [34], researchers start to integrate the semantics from these foundation models into their reconstruction. SA3D [7] is the pioneering method that extends 2D SAM masks across images into 3D consistent object masks. GARField [19] introduces hierarchical grouping, leveraging the physical scales of 2D masks to group objects of different sizes and represent scenes at different granularity scales. Gaussian Grouping [57] is a pioneering Gaussian-based approach that utilizes consistent object IDs generated by video tracker DEVA [8] to render object IDs into camera views. Concurrent work such as SAGA [6] utilizes the masks generated from SAM, combined with scales obtained from 3D data inspired by GARField, to segment objects of various sizes. Notably, these methods are designed for static scenes. To date, only two concurrent work, SA4D [16] and DGD [25], addresses the segmentation task in dynamic 3DGS. DGD extends the rasterization pipeline of 3DGS to be able to render semantic feature of each Gaussian. The rendered Gaussian feature map is then trained with the supervision of DINOv2 [34] or CLIP [42] to mimic the output from their feature extractors. The biggest drawback of such supervision is anticipated longer training times as the dimensions of the features generated by DINOv2 and CLIP are quite large (384 and 512). Similar to Gaussian Grouping and its derivatives [11, 31], SA4D relies on video trackers to associate masks across different views to produce consistent object IDs. These approaches face significant limitations, as the mask association often fails to offer consistent object IDs in dynamic scenes, such as the oven-mitts sequence from HyperNeRF [36] as shown in Fig. 2.
2.2 Scene Editing Applications
Learning semantically-aware latent space opens opportunities for interactive downstream applications, allowing users to manipulate scenes on an object level. Several works proposed editing techniques for radiance field representations [13, 49], which are limited to a single object. On the other hand, scene-based methods [53, 24] learn a separate radiance field per object and render the full scene by combining learned opacities along the viewing rays. While the structure is modular, its editing application is limited to a pre-defined number of objects. Our method is guided by user interaction and is capable of operating on any object in the scene. Another line of work distills features such as DINO [34] into volumetric representation and shows editing capabilities on learned continuous feature field [48]. Due to the implicit representation, scene editing requires a re-assessment of the final consistency by rendering from novel scenes, which impedes the model’s applicability to real-time interactive applications.
At the same time, Gaussian splatting methods mitigate the runtime limitation of NeRFs with their explicit point-based representation. SAGA [57] proposes Local Gaussian Editing scheme, which localize the Gaussians and manipulate them according to the editing task. Feature3DGS [60] distills features from 2D foundation models such as SAM [43] and LSeg [26] enabling promptable and language-driven 3D edits. Nonetheless, these methods operate only on static scenes, while our method is capable of editing dynamic objects and ensuring temporal consistency of the edits. DGD [25] shows semantic texture editing performance by utilizing a pre-trained diffusion model and fine-tuning dynamic 3DGS representation with the reconstruction loss. On the other hand, we demonstrate a variety of editing tasks, including style transfer, object removal, and composition. Despite offering several scene editing opportunities, SA4D [16] has limited user interactivity and requires prior knowledge of object identities. Our framework contains a user-friendly graphical interface that allows operation with a simple text prompt or a mouse click.
3 Method
In this section, we introduce our novel framework SADG. As illustrated in Fig. 1 the pipeline comprises two main components: dynamic geometry reconstruction (Sec. 3.1) and Gaussian feature learning (Sec. 3.2). We adopt the Deformable-3DGS [56] pipeline to reconstruct the 4D scene. Once the 4D reconstruction is learned, we freeze this representation and proceed with Gaussian feature learning using SAM [20] masks and our novel contrastive learning objective in the rendered feature space.
3.1 Dynamic Geometry Reconstruction
We follow the approach proposed in [56] and describe it briefly for completeness. Unlike Dynamic3DGS [30] model, where the reconstruction is stored per frame, resulting in high memory consumption, an MLP learns per-Gaussian deformation with respect to the static canonical space , which is defined as
is the center position of the i-th Gaussian, is the quaternion representing the rotation, is the scaling vector, refers to the opacity, and is the Spherical Harmonic Coefficients encoding the color information () as per [56].
The resulting Gaussians at timestamp are defined as
Once is obtained, we proceed with the standard rasterization pipeline [17] to render the image . The color loss is computed with the ground truth image as
where refers to the weighting parameter for the structural similarity loss term.
3.2 Gaussian Feature Learning
Problem statement Most of the existing works [16, 57] utilize a video tracker to provide consistent object mask IDs across views and render the segmentation field directly by predicting the per-pixel probability of different object classes. However, such approaches have a significant bottleneck, namely their reliance on perfect object mask association across views. When mask inconsistencies of the same object occur between views, the entire optimization pipeline can break down, leading to sub-optimal results or failure.
Therefore, we propose a method that aims to learn the 3D semantics of objects and leverage them directly for downstream tasks across different views. In other words, instead of predicting per-pixel probabilities during rendering, we focus on integrating semantic information into Gaussians and leverage semantic constancy for learning multi-view consistent spatio-temporal feature field.
Representation Given 4D reconstruction, we incorporate semantic information into the scene. Specifically, we extend its -th Gaussian with a 32-dimensional semantic feature vector , the so-called “Gaussian feature”. The semantically-aware feature field of a scene represented with Gaussians can be expressed as
Rendering Gaussian Features The learned Gaussian features can be rendered similarly to [17, 60] with point-based -rendering [23, 22]. Given a sorted subset of Gaussians that are related to pixel , the resulting rendered features can be computed as follows:
Existing methods [6] use smooth Gaussian features instead, as they help to filter out noisy 3D Gaussian features with high similarity scores corresponding to unrelated objects. The smooth Gaussian feature for the -th Gaussian is computed as per Eq. 6, where denotes a set of nearest neighbors [9] of the -th Gaussian based on its 3D position. We adopt this convention in our work and utilize smooth feature vectors in the rendering process (Eq. 5).
Our contrastive semantically-aware learning objective
The generated set of SAM masks comprise binary masks of a given image , where each mask corresponds to a different object. In practice, we do not consider all masks from SAM to alleviate memory and computational requirements. Instead, we randomly sample masks for further usage.
We define the pixel-mask correspondence vector in Eq. 7, which captures mask information for a certain pixel coordinate (, ) across all binary masks in . This vector is similar to a one-hot encoding technique; however, there is no guarantee that a pixel belongs to a single mask due to 2D segmentation inaccuracies. We visually demonstrate the formulation of in Fig. 4.
where and .
In practice, we only sample pixels from an image such that . Mathematically, we define the sampled pixel coordinates as a set .
The pairwise similarities of the sampled pixel-mask correspondence vectors can be retrieved by computing the Gram matrix of their stacked matrix . All entries of the resulting Gram matrix are then bounded to either 0 (negative pair) or 1 (positive pair) to form a mask-based correspondence matrix among sampled pixels , formally defined as
Additionally, we formulate a correspondence matrix for the sampled rendered features . Analogous to the mask-based correspondence matrix , we compute a feature-based Gram matrix, capturing the pairwise similarities between the features as follows:
where is defined as follows:
Finally, we have all the necessary ingredients to formulate the contrastive learning objective based on a hard positive and negative mining scheme. Guided by the mask-based correspondence matrix, we enforce features to be close for pixels from the same object mask and far apart for the distinct objects. Furthermore, we concentrate on hard pixel pairs. Specifically, we sample positive samples with low similarity scores and negative samples with high similarity scores. With this, we formally define the optimization objective as
Note that is an indicator function. It returns 1 when , and 0 otherwise. stands for hard positive which is defined as and refers to a hard negative which is defined as . Here, and are the positive and negative thresholds for pairwise similarity respectively. As both and are symmetric matrices following the property of the Gram matrix, we only consider entries in the upper triangular part and exclude the diagonal to avoid repeated computation. We show rendered Gaussian features learned with our contrastive scheme in Fig. 1. For demonstration, they are compressed into via PCA [33], allowing each rendered feature to be represented with RGB color.
Gaussian Feature Clustering To enable interactive applications with the learned 4D representation, we deploy the DBSCAN algorithm [12] to cluster the Gaussian features into several groups and generate the so-called Gaussian Clusters as defined in Eq. 12. provides the corresponding cluster ID for -th Gaussian feature:
Note that refers to the total number of clusters. We demonstrate them visually in Fig. 1. As can be observed, they represent distinct objects in 4D space.
Since a typical scene contains a large number of Gaussians (more than 300k), we adopt the sub-sampling scheme proposed by [6], where only 2% of the Gaussian features are used for clustering. We propagate the cluster IDs to the remaining of the Gaussians by computing the feature similarity between cluster-representative Gaussians. In the supplementary material, we demonstrate that the performance is not affected by a significant reduction in the number of Gaussians contributing to the generation of clusters.
4 Experimental Results
We evaluate the performance of SADG across several well-known datasets to validate its effectiveness. We begin by introducing the datasets in Sec. 4.1. Since there are no established benchmarks for validating the segmentation performance on novel views in dynamic scenes, we provide details on how we extend these different datasets with segmentation benchmarks (Sec. 4.2). Next, we explore the implementation details and various configurations used to train and test methods on these benchmarks (Sec. 4.3). We present segmentation results of our model, along with both quantitative and qualitative comparisons to other related works [6, 57, 25] in Sec. 4.5. Last but not least, we demonstrate capabilities of our learned semantic representation on a number of downstream tasks including scene editing and selection in our a graphical user interface (Sec. 4.6).
4.1 Datasets
NeRF-DS-Mask | HyperNeRF-Mask | Neu3D-Mask | Immersive-Mask | Technicolor-Mask | ||||||
Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
Gaussian Grouping [57] | 0.8351 | 0.9729 | 0.8341 | 0.9771 | 0.8864* | 0.9932* | 0.7673* | 0.9776* | 0.7979* | 0.9800* |
SAGA [6] | 0.8072 | 0.9644 | 0.7660 | 0.9579 | 0.6941 | 0.9516 | 0.7395 | 0.9747 | 0.7913 | 0.9792 |
DGD [25] | 0.7125 | 0.9551 | 0.8297 | 0.9777 | 0.7721 | 0.9851 | 0.7981 | 0.9850 | 0.7791 | 0.9728 |
SADG (Ours) | 0.8719 | 0.9826 | 0.8663 | 0.9845 | 0.9022 | 0.9945 | 0.9234 | 0.9945 | 0.9308 | 0.9917 |
NeRF-DS dataset [52] contains scenes with challenging transparent and shiny moving objects recorded by two monocular cameras. Five sequences are used for reconstruction and evaluation including as_novel_view, basin_novel_view, cup_novel_view, press_novel_view, and plate_novel_view.
HyperNeRF dataset [36] is recorded using a single monocular camera . The camera moves while capturing the dynamic scenes involving different human-object interactions. In our experiments, we use ten sequences: americano, split-cookie, oven-mitts, espresso, chickchicken, hand1-dense-v2, torchocolate, slice-banana, keyboard, and cut-lemon1.
Neu3D dataset [27] contains videos captured by a multi-view rig with to cameras each at 30 FPS. The scenes involve a person performing cooking tasks in the kitchen. In total, five scenes are used in our experiments: coffee_martini, cook_spinach, cut_roasted_beef, flame_steak, and sear_steak.
Google Immersive dataset [2] consists of videos captured by a 46-camera rig at 30 FPS. The cameras are fish-eye cameras and are distributed on the surface of a hemispherical, 92cm diameter dome. Four scenes are used in our experiments: 01_Welder, 02_Flames, 10_Alexa_1, and 11_Alexa_2.
Technicolor Light Field dataset [44] captures scenes using a camera array at 30 FPS. We use the undistorted images provided by the authors, and four scenes are used in our experiments: Birthday, Fabien, Painter, and Theater.
4.2 Segmentation Benchmarks
To date, while numerous benchmarks exist for dynamic video segmentation, such as DAVIS [38, 40, 3, 4] and VOS [51, 54, 55], there are still no well-established segmentation benchmarks for dynamic novel view synthesis.
Such benchmarks are crucial for evaluating a comprehensive understanding of scenes, both geometrically and semantically. Existing datasets, such as Replica [47], ScanNet [10], and ScanNet++ [58], include ground-truth labels for evaluating panoptic segmentation performance, but they do not cover dynamic sequences. To address this gap, we manually annotate dynamic sequences in our experiments and propose the NeRF-DS-Mask, HyperNeRF-Mask, Neu3D-Mask, Immersive-Mask, and Technicolor-Mask benchmarks. Utilizing the powerful SAM2 model [43], which provides consistent object masks across video sequences, we manually define the objects of interest for each sequence to evaluate their segmentation performance in our model. Examples of these mask objects can be seen in the supplementary.
4.3 Implementation Details
We employ SAM [20] to generate the anything-masks for Gaussian feature learning. However, the resulting masks are memory-inefficient because they are stored in a dense format () with a lot of pixels with zero values. To optimize memory usage, we save the masks as a bit array.
Given a permuted tensor version of an RGB image , SAM generates masks . We first flatten , then convert flattened ()-dimensional array into a bit-array. The final output is stored in a dictionary with the required information to re-create the original mask tensor: a number of masks , image dimensions , and the compressed bit array, i.e., . This approach significantly reduces the storage used and is particularly useful for multi-view datasets, where numerous masks from different views are generated. For instance, a raw mask generated from an image in coffee_martini occupies around 80 MB, but the dictionary representation is significantly reduced to approximately 1 MB. Due to space limits, we provide training details for different datasets in the supplementary.
4.4 Evaluation Protocol and Baselines
Given ground truth masks we use Mean Intersection over Union (mIoU) and Mean Pixel Accuracy (mAcc) to quantitatively evaluate segmentation performance of the models. We compare our model with SAGA [6], Gaussian Grouping [57] and DGD [25]. Note that both SAGA and Gaussian Grouping are designed for static scenes and do not include a module to encode temporal information. To address this limitation, we integrate our 4D reconstruction backbone into their pipeline, extending their approach to effectively handle dynamic scenes. Moreover, Gaussian Grouping needs the precomputed object IDs from DEVA [8], whose labels are inconsistent for multi-view camera setup. Thus, we train their semantic module on a monocular stream from the camera closest to the test view. Furthermore, all reported baselines utilize the same pre-trained 4D reconstruction, with semantics learned according to each pipeline. This allows us to isolate and evaluate the performance of each model based on its semantic learning capabilities. After training, we manually perform click prompts in the rendered novel views of each scene, which are trained by different approaches to select the objects of interest defined in the benchmark. Apart from the quantitative results, we also present qualitative results from various benchmarks, as certain insights cannot be fully captured through numerical evaluations alone. Since the official source code for SA4D [16] is not available, we report the comparison based on unofficial implementation in the supplementary.
4.5 Segmentation Results
As shown in Tab. 1, SADG consistently outperforms all other methods across all benchmarks in terms of average mIoU and mAcc. We also select some sequences to visually illustrate the quality of the segmentation of each model in Fig. 5. These comparisons reveal that SADG qualitatively surpasses existing models and exhibits multi-view consistent segmentation performance. Notably, SADG maintains strong generalization to test camera views in multi-view sequences such as flame_steak, Painter, and 11_Alexa_2. This robust performance can be attributed to the compact Gaussian features learned through the proposed Gaussian feature training. In contrast, DGD segments objects by comparing cosine similarity scores between features retrieved via click prompts. However, determining a suitable threshold that generalizes across various scenes is neither straightforward nor intuitive.
SADG employs clustering based on the Gaussian features to group the Gaussians into several segments. As this operation is done in 3D space, SADG achieves cross-view consistency and does not need to associate objects from different views. This advantage is evident in cut-lemon1 example in Fig. 5, where wrongly associated object IDs can cause segmentation failure for Gaussian Grouping.
4.6 Scene Editing Applications
In addition to accurate segmentation capabilities and multi-view consistent rendered masks, we demonstrate the effectiveness of our learned semantic feature field in a range of downstream tasks. Specifically, we focus on interactive editing applications and develop a graphical user interface (GUI) to grant full user control over the representation. Thanks its explicit nature and our semantic abstraction to Gaussian clusters, we achieve real-time operation capability in handling many editing tasks. We provide snapshots of the GUI in the supplementary.
Besides simple click prompts in the novel view of dynamic scenes for object selection, we also support text-based prompts to make object selection more intuitive. To enhance 3D content editing in dynamic scenes, we integrate style transfer capabilities for the selected objects. By segmenting objects directly in 3D, we facilitate scene composition and object removal through straightforward manipulation of the selected Gaussians. Due to space limit, we shall provide more qualitative results in the supplementary. In the following we describe the implementation details of editing tasks based on our semantically-aware feature field.
Text Prompt. Inspired by Gaussian Grouping [57], we employ Grounding DINO [28] to generate 2D mask based on the text prompt in the novel view. The 2D mask is reprojected into the 3D space with the rendered depth information. Finally, we associate each reprojected point with its nearest Gaussian in the 3D scene and retrieve its corresponding cluster. If the number of associated Gaussian of a given cluster is more than a threshold, the cluster would be selected. The example of text prompts on different sequences is shown in Fig. 6 (a).
Scene Composition. We also perform qualitative results on scene composition. As the segmentation is performed in 3D, we can simply put the selected Gaussians from the dynamic scene into another static or dynamic scene. The scales and coordinate transformation need to be adapted. We show some examples on scene composition in Fig. 6.
Object Style Transfer. We adopt the static Gaussian style transfer from StyleSplat [14] for dynamic scenes. Given a rendered image and a style image , we extract their feature maps (, respectively) from VGG16 [46] and optimize only the SHs of the Gaussians of the selected cluster with the nearest-neighbor feature matching (NNFM) loss [59]. Fig. 6 (c) illustrates an example of style transfer on the segmented object.
Object Removal. We can also remove object by simply deleting Gaussians belonging to the specific cluster as illustrated in Fig. 6 (d). It is worth noting that no inpainting is required due to the underlying 3D representation, where exposed parts are learned by multi-view observations.
5 Conclusion
We introduced a novel framework for dynamic scene understanding, which enables multi-view consistent segmentation without any object tracking supervision. Our SADG effectively combines dynamic 3D Gaussian Splatting 3DGS [56] and 2D SAM [20] masks in a contrastive learning objective, which lifts semantic information into 3D space and learns expressive Gaussian features based on hard pixel cases. This results in cross-view consistency when rendering segmented objects, enhancing the quality and coherence of object segmentation across different views. Evaluated on various novel-view datasets, SADG shows superior performance both quantitatively and qualitatively. We further demonstrated the effectiveness of the learned feature field on downstream editing tasks such as point and text prompts, style transfer, object removal, and scene composition. Our approach sets a strong foundation for further research into dynamic scene understanding and scene editing, especially in complex multiview scenarios.
Supplementary Material
In the supplementary, we provide further implementation details in Sec. 6 and Sec. 7. A range of ablation studies on different components of our pipeline is given in Sec. 8. We compare our SADG with the concurrent work on 4D segmentation and editing, Segment Any 4D Gaussians (SA4D), in Sec. 9. Several examples from our segmentation benchmark are given in Sec. 11. We present detailed quantitative results on all benchmarks in Sec. 10. We finally provide qualitative results of our method on segmentation (Sec. 12, Sec. 14) and editing (Sec. 13).
6 Training Implementation Details
We implement the whole pipeline in PyTorch [37]. We reuse the codebase from Deformable-3DGS [56] and extend the differential Gaussian rasterization pipeline to render the Gaussian features similar to [57]. In the following, we introduce the detailed training setup for different datasets.
6.1 NeRF-DS and HyperNeRF
For the NeRF-DS sequences, we use the left camera for training and the right camera for testing. In the HyperNeRF sequences, training frames are sampled every 4 frames, starting from the first frame, while testing frames are selected every 4 frames, starting from the third frame. This follows the interp-sequence data-loading pipeline from the original HyperNeRF paper [36], ensuring that all testing frames are unseen during training. Note that we use -downsized images for all experiments.
NeRF-DS-Mask | HyperNeRF-Mask | Neu3D-Mask | Immersive-Mask | Technicolor-Mask | ||||||
Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
SADG (w/o filtering) | 0.8719 | 0.9826 | 0.8368 | 0.9790 | 0.8738 | 0.9923 | 0.9154 | 0.9942 | 0.9258 | 0.9904 |
Ours (w/ filtering) | 0.8719 | 0.9826 | 0.8663 | 0.9845 | 0.9022 | 0.9945 | 0.9234 | 0.9945 | 0.9308 | 0.9917 |
NeRF-DS-Mask | HyperNeRF-Mask | Neu3D-Mask | Immersive-Mask | Technicolor-Mask | ||||||
Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
SADG (K=0) | 0.8228 | 0.9706 | 0.8360 | 0.9787 | 0.8661 | 0.9920 | 0.8725 | 0.9909 | 0.9259 | 0.9908 |
Ours (K=16) | 0.8719 | 0.9826 | 0.8368 | 0.9790 | 0.8738 | 0.9923 | 0.9154 | 0.9942 | 0.9258 | 0.9904 |
SADG (K=32) | 0.8754 | 0.9828 | 0.8331 | 0.9780 | 0.8702 | 0.9922 | 0.9217 | 0.9944 | 0.9255 | 0.9904 |
Dynamic Geometry Reconstruction. We adopt the default hyper-parameters from Deformable-3DGS [56]. Initially, the Gaussians are warmed up without the deformation MLP for the first 3k iterations to better capture the positions and shape stably as suggested in [56], after which the MLP optimization begins, and the 3D Gaussian and the deformation MLP are trained jointly. The densification of Gaussians for adaptive density control introduced in 3DGS [17] is set to run until 15k iterations. The dynamic geometry reconstruction is trained for a total of 20k iterations.
Gaussian Feature Learning. After the dynamic reconstruction is completed, we freeze model’s weights and begin Gaussian feature learning for an additional 10k iterations. For each image, we randomly sample 25 masks () generated by SAM [20] and 5k pixels (), with K = 16 to get smooth Gaussian features.
6.2 Neu3D
Each video contains 300 frames, and we begin by down-scaling the images by a factor of 2, resulting in a resolution of . The center camera (camera_00) is used for testing, while the remaining cameras are for training. Initially, we run SfM [45] across all the first frames, excluding camera_00, to get the initial point cloud.
Dynamic Geometry Reconstruction. We follow the same pipeline for dynamic geometry reconstruction as described in Tab. 3, except the densification of Gaussians is set to 8k iterations.
Gaussian Feature Learning. After the dynamic reconstruction is completed, we freeze the model’s weights and begin Gaussian feature learning for an additional 10k iterations. For each image, we randomly sample 50 masks () generated by SAM [20] and 10k pixels (), with K = 16 to get the smooth Gaussian feature.
6.3 Google Immersive
Each video contains frames captured by fish-eye cameras with 300 frames. We first undistort the images and align the principal points to the center, then downscale them to a resolution of .
NeRF-DS-Mask | HyperNeRF-Mask | Neu3D-Mask | Immersive-Mask | Technicolor-Mask | |||||||||||
Avg. Num. Gaussians | 186455.4 | 565614.7 | 678566 | 985713 | 457677.75 | ||||||||||
Method | mIoU | mAcc | time | mIoU | mAcc | time | mIoU | mAcc | time | mIoU | mAcc | time | mIoU | mAcc | time |
Ours (2%) | 0.8719 | 0.9826 | 0.4068 | 0.8368 | 0.9790 | 2.3117 | 0.8738 | 0.9923 | 2.2506 | 0.9154 | 0.9942 | 9.6083 | 0.9258 | 0.9904 | 1.6118 |
SADG (10%) | 0.8707 | 0.9828 | 1.5782 | 0.8384 | 0.9794 | 11.8430 | 0.8792 | 0.9930 | 6.1878 | 0.9160 | 0.9939 | 71.2429 | 0.9283 | 0.9908 | 6.8211 |
SADG (25%) | 0.8762 | 0.9832 | 3.9425 | 0.8461 | 0.9809 | 43.2682 | 0.8886 | 0.9935 | 18.8976 | 0.9180 | 0.9943 | 221.7642 | 0.9297 | 0.9907 | 19.0767 |
Similar to Neu3D, the center camera (camera_0001) is used for testing, while the other cameras are used for training. We generate the initial point cloud using SfM across all the first frames, excluding the test camera, and apply the same hyper-parameters as in Neu3D in Sec. 6.2.
6.4 Technicolor Light Field
We use the undistorted images provided by the authors, aligning the principal points to the center and resizing the images to a resolution of . Similar to Neu3D and Google Immersive dataset, the center camera (camera_0000) is used for testing, while the other cameras are used for training. We generate the initial point cloud using SfM across all the first frames excluding the test camera and apply the same hyper-parameters as in Neu3D in Sec. 6.2.
7 Testing Implementation Details
After the dynamic geometry reconstruction and the Gaussian features are properly trained, we proceed to our GUI to select the object of interest and evaluate the performance of the model by mIoU and mAcc as discussed in Sec. 4.4.
Given a clicked point on the novel view, we first use the known camera intrinsics to compute the inverse intrinsic matrix . With the depth information rendered by the dynamic reconstruction of the Gaussians at time , we can reproject the 2D clicked point to 3D coordinates in the camera frame. Finally, using the known camera extrinsics (rotation matrix and translation vector ) for the world-to-camera transformation, we convert the points back into the world coordinate frame. We can then use this point coordinate to query its nearest neighbor Gaussian using KNN. The cluster ID corresponding to its nearest neighbor Gaussian is subsequently retrieved.
We provide interactive illustrations in our GUI to segment target objects as follows:
1.
Open the GUI and run DBSCAN (Click button ”Clustering”) to group Gaussians into different clusters based on the learned Gaussian features as illustrated in Fig. 7.
2.
Click the objects of interest in the novel view to segment the target objects as shown in Fig. 8.
3.
Click the button ”Render Object” to render the selected object to test views for evaluating the performance.
HyperNeRF-Mask | Neu3D-Mask | Immersive-Mask | Technicolor-Mask | |||||
Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
SA4D [16] | 0.7767 | 0.9633 | 0.8832* | 0.9391* | 0.7987* | 0.9835* | 0.8271* | 0.9734* |
SADG (Ours) | 0.8663 | 0.9845 | 0.9022 | 0.9945 | 0.9234 | 0.9945 | 0.9308 | 0.9917 |
8 Ablation Studies
8.1 Filtering
We noticed that the raw segmented object from the selected Gaussian clusters sometimes has some artifacts near the outline of the objects as illustrated in Fig. 9. Therefore, we adopt a simple yet effective filtering step to filter out those artifacts. The procedure of the filtering is as follows:
In each selected Gaussian cluster , we compute the mean Gaussian features in the selected cluster .
We compute the per-Gaussian similarity score between the Gaussian features inside the selected cluster and mean Gaussian features of the selected cluster with cosine similarity.
Delete Gaussians that have lower similarity scores than the threshold .
As shown in Tab. 2, the proposed filtering step can effectively improve the average mIoU and mAcc for each benchmark.
8.2 Smooth Gaussian Features
In Sec. 3.2 we introduced smooth Gaussian features (Eq. 6). We provide a detailed analysis of the selection of the number of nearest neighbors for the KNN algorithm in Tab. 3. Overall, when smooth Gaussian features are enabled ( and ), the performance is better than when smooth Gaussian features are disabled (). Since there is no significant improvement between and , we eventually chose for our pipeline to avoid additional overheads with the KNN sampling.
8.3 Gaussian Feature Clustering
In Sec. 3.2 we additionally introduced Gaussian Feature Clustering and the sub-sampling strategy to run the clustering in real-time with scenes that normally contain more than 300k Gaussians. We perform time analysis with different sub-sampling numbers of Gaussians (2%, 10%, and 25%) on RTX3080 and Intel i7-12700. For each scene, we compute the time for clustering using Python timeit package and take the average clustering time for five runs per scene. We report the average clustering time across all scenes in each benchmark and the corresponding mIoU and mAcc for our object of interest in Tab. 4. Overall, with different sub-sampling of Gaussians (2%, 10%, and 25%), the mIoU and mAcc stay almost unchanged. However, as increasing the number Gaussians results in a longer clustering time, we believe that 2% is the optimal sub-sampling number for clustering.
8.4 DBSCAN vs. K-Means
There are various clustering techniques and K-means [29, 32, 15, 39] is one of the most widely used. K-means partitions data into clusters, where each feature belongs to the cluster with the nearest centroid (mean of the cluster). The value must be predefined, and the algorithm iteratively adjusts the centroids to minimize the sum of squared distances between features and their corresponding cluster centers. Therefore, the performance of the algorithm heavily depends on the optimal value of .
To investigate this, we conducted an additional ablation study comparing the rendered segmentation produced by K-means with different values of and DBSCAN. As shown in Fig. 10, the choice of significantly impacts the result of rendered segmentation. Since different sequences contain varying numbers of objects, we believe that employing a method with an adaptive number of clusters, such as DBSCAN, validates our design choices with additional flexibility and no loss on the segmentation accuracy.
8.5 Time and Storage Analysis
We performed a comprehensive time analysis of SADG using an RTX 3080 GPU and an Intel i7-12700 CPU on the NeRF-DS [52] dataset. To evaluate its performance, we compared it with DGD [25], the only publicly available model addressing segmentation in dynamic novel views. The results are summarized in Tab. 6.
Most existing approaches for novel view segmentation [57, 6, 16] require precomputing SAM masks or object masks prior to training. In contrast, DGD dynamically computes DINOv2 feature maps for supervision during training. For a fair comparison, we precomputed the DINOv2 feature maps for all training images in DGD and stored them beforehand on an SSD, alongside our precomputed SAM masks.
As shown in Tab. 6, DGD [25] spends little preprocessing time for DINOv2 features (approximately 5 seconds for all training images per sequence). Although our SADG requires preprocessing time to generate SAM masks, this process only needs to be done once before training. DGD assigns a 384-dimensional semantic feature to each Gaussian. Rendering these high-dimensional features into 2D is computationally inefficient, significantly slowing down the training process. SADG uses 32-dimensional Gaussian features, which can be rendered much more efficiently. As a result, SADG achieves over 3 faster total time (preprocessing + training) than DGD while only requiring 4.5 less space in the storage.
In addition, we evaluated the processing time for click prompts in our GUI. For smaller scenes with less than 200k Gaussians such as as_novel_view, segmenting an object via a click prompt takes approximately 0.05 seconds. For larger scenes with more than 1,000k Gaussians captured by a multi-view camera setup, such as 01_Welder, the processing time per click prompt is around 0.4 seconds.
Avg. Preprocessing Time (min) | Avg. Training Time (min) | Total Time (min) | Avg. Size of Gaussians (MB) | |
DGD [25] | 0.08 | 103 | 103 | 285 |
SADG (Ours) | 16 | 16 | 32 | 60 |
9 Comparison Against Segment Any 4D Gaussians Model
We provide additional quantitative and qualitative comparisons with the concurrent work on 4D segmentation and editing, SA4D [16]. The pipeline consists of three stages: (1) dynamic 3D reconstruction based on 4D-GS model [50], (2) identity feature field learning, and (3) post-processing with outlier removal. Since the authors have not released the official code, we use the unclean version from https://github.com/guanjunwu/sa4d for training and testing the model on our benchmark. We do not provide the performance on NeRf-DS-Mask with specular and transparent moving objects in Tab. 5 since 4D-GS fails to converge on most of the sequences apart from two (basin_novel_view, press_novel_view). Similarly, we exclude two sequences from the HyperNeRF-Mask dataset, such as hand1-dense-v2 and espresso due to 4D reconstruction failure.
In quantitative assessment, we select the top 30 segmentation results to more closely reflect the test set used in the original SA4D benchmark (not released). We also uniformly sub-sample point cloud (100K points) from SfM [45]. As seen in Tab. 5, our SADG significantly outperforms SA4D on all datasets. From the qualitative results in Fig. 12 and Fig. 13, we notice that rendered masks are larger than the underlying object and include a lot of unrelated regions even after the outlier removal in the post-processing step, which negatively affects objects’ segmentation accuracy. We further demonstrate some editing comparisons in Fig. 17, Fig. 16, and Fig. 15. Due to inaccurate segmentation, composition, and removal, SA4D results contain many spurious regions, degrading the editing quality.
10 Quantitative Results per Sequence
For completeness to the average numbers per datasets (Tab. 1), we report the performance of our method individually on all benchmarks in Tab. 7, Tab. 8, Tab. 9, Tab. 10, and Tab. 11.
11 Segmentation Benchmarks
12 Additional Qualitative Results
We provide additional qualitative results in Fig. 12 and Fig. 13 to further showcase the effectiveness of SADG over other methods. While SAGA [7] suffers from degenerate Gaussians, our reconstruction backbone and effective post-processing mitigate the issue. At the same time, rendered masks from other methods, such as DGD [25] and SA4D [16], are inaccurate and introduce spurious regions. Our method demonstrates the best performance among all approaches.
13 Further Scene Editing Examples
We present further scene composite examples of SADG in Fig. 14. We also compare the quality of object removal and scene composition with SA4D [16] in Fig. 16 and Fig. 17. Overall, as discussed in Sec. 9, SADG outperforms SA4D both quantitatively and qualitatively.
as_novel_view | basin_novel_view | cup_novel_view | press_novel_view | plate_novel_view | Average | |||||||
Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
Gaussian Grouping [57] | 0.8140 | 0.9487 | 0.8083 | 0.9612 | 0.9020 | 0.9936 | 0.7803 | 0.9912 | 0.8710 | 0.9696 | 0.8351 | 0.9729 |
SAGA [6] | 0.8235 | 0.9542 | 0.6437 | 0.9140 | 0.8833 | 0.9922 | 0.8119 | 0.9918 | 0.8736 | 0.9699 | 0.8072 | 0.9644 |
DGD [25] | 0.7135 | 0.9266 | 0.6685 | 0.9291 | 0.694 | 0.9752 | 0.6457 | 0.9822 | 0.8406 | 0.9625 | 0.7125 | 0.9551 |
SADG (Ours) | 0.9162 | 0.9799 | 0.8740 | 0.9785 | 0.8864 | 0.9924 | 0.8098 | 0.9916 | 0.8733 | 0.9707 | 0.8719 | 0.9826 |
americano | chickchicken | cut-lemon1 | espresso | hand | keyboard | |||||||
Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
Gaussian Grouping [57] | 0.8980 | 0.9969 | 0.9436 | 0.9867 | 0.7153 | 0.9394 | 0.4968 | 0.9642 | 0.8865 | 0.9814 | 0.9159 | 0.9861 |
SAGA [6] | 0.7325 | 0.9873 | 0.8633 | 0.9648 | 0.5263 | 0.8398 | 0.6395 | 0.9798 | 0.9088 | 0.9859 | 0.8070 | 0.9645 |
DGD [25] | 0.7748 | 0.9912 | 0.8317 | 0.9589 | 0.8583 | 0.9718 | 0.7240 | 0.9871 | 0.8936 | 0.9838 | 0.8451 | 0.9733 |
SADG (Ours) | 0.8144 | 0.9922 | 0.9308 | 0.9835 | 0.8795 | 0.9769 | 0.7164 | 0.986 | 0.9006 | 0.9849 | 0.8952 | 0.9827 |
oven-mitts | slice-banana | split-cookie | torchocolate | Average | ||||||||
Gaussian Grouping [57] | 0.7420 | 0.9505 | 0.9095 | 0.9721 | 0.9134 | 0.9960 | 0.9196 | 0.9981 | 0.8341 | 0.9771 | ||
SAGA [6] | 0.8667 | 0.9766 | 0.7323 | 0.8960 | 0.8186 | 0.9906 | 0.7652 | 0.9933 | 0.7660 | 0.9579 | ||
DGD [25] | 0.8595 | 0.9755 | 0.8408 | 0.947 | 0.8467 | 0.9925 | 0.8223 | 0.9954 | 0.8297 | 0.9777 | ||
SADG (Ours) | 0.9295 | 0.9876 | 0.8782 | 0.9617 | 0.8581 | 0.9930 | 0.8599 | 0.9964 | 0.8663 | 0.9845 |
coffee_martini | cook_spinach | cut_roasted_beef | flame_steak | sear_steak | Average | |||||||
Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
Gaussian Grouping [57] | 0.8295 | 0.9890 | 0.8974 | 0.9941 | 0.9512 | 0.9972 | 0.8279 | 0.9900 | 0.9261 | 0.9959 | 0.8864 | 0.9932 |
SAGA [6] | 0.2201 | 0.8081 | 0.8125 | 0.9881 | 0.6982 | 0.9767 | 0.8439 | 0.9911 | 0.8959 | 0.9941 | 0.6941 | 0.9516 |
DGD [25] | 0.7875 | 0.9865 | 0.815 | 0.9883 | 0.817 | 0.9877 | 0.6771 | 0.9776 | 0.7638 | 0.9854 | 0.7721 | 0.9851 |
SADG (Ours) | 0.9120 | 0.9948 | 0.9129 | 0.9951 | 0.9103 | 0.9947 | 0.8716 | 0.9930 | 0.9044 | 0.9947 | 0.9022 | 0.9945 |
01_Welder | 02_Flames | 10_Alexa_1 | 11_Alexa_2 | Average | ||||||
Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
Gaussian Grouping [57] | 0.7920 | 0.9736 | 0.8364 | 0.9936 | 0.7391 | 0.9840 | 0.7018 | 0.9590 | 0.7673 | 0.9776 |
SAGA [6] | 0.8062 | 0.9737 | 0.8463 | 0.9940 | 0.4794 | 0.9513 | 0.8260 | 0.9797 | 0.7395 | 0.9747 |
DGD [25] | 0.873 | 0.9854 | 0.7775 | 0.9914 | 0.8107 | 0.9899 | 0.7312 | 0.9734 | 0.7981 | 0.9850 |
SADG (Ours) | 0.8975 | 0.9877 | 0.8808 | 0.9955 | 0.9367 | 0.9970 | 0.9786 | 0.9979 | 0.9234 | 0.9945 |
Birthday | Fabien | Painter | Theater | Average | ||||||
Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
Gaussian Grouping [57] | 0.9324 | 0.9908 | 0.9406 | 0.9758 | 0.8759 | 0.9888 | 0.4428 | 0.9646 | 0.7979 | 0.9800 |
SAGA [6] | 0.8541 | 0.9779 | 0.9371 | 0.9943 | 0.9670 | 0.9867 | 0.4071 | 0.9579 | 0.7913 | 0.9792 |
DGD [25] | 0.7203 | 0.9573 | 0.9177 | 0.9667 | 0.7959 | 0.9788 | 0.6826 | 0.9885 | 0.7791 | 0.9728 |
SADG (Ours) | 0.8839 | 0.9833 | 0.9734 | 0.9889 | 0.9719 | 0.9976 | 0.8941 | 0.9969 | 0.9308 | 0.9917 |
14 Visualization of the Rendered View, Rendered Segmentation, Segmented Object
We also provide the visualizations of the rendered novel view, the rendered segmentation, and the segmented objects in Fig. 18, Fig. 19, and Fig. 20. The rendered segmentation of SADG is based on Gaussian clustering, which is performed in 3D space, resulting in consistent segmentation in the dynamic scene.

### v2

@@
-SADG: Segment Any Dynamic Gaussian Without Object Trackers
+TRASE: Tracking-free 4D Segmentation and Editing
 Abstract
-Understanding dynamic 3D scenes is fundamental for various applications, including extended reality (XR) and autonomous driving.
-Effectively integrating semantic information into 3D reconstruction enables holistic representation that opens opportunities for immersive and interactive applications.
-We introduce SADG, Segment Any Dynamic Gaussian Without Object Trackers, a novel approach that combines dynamic Gaussian Splatting representation and semantic information without reliance on object IDs.
-In contrast to existing works, we do not rely on supervision based on object identities to enable consistent segmentation of dynamic 3D objects.
-To this end, we propose to learn semantically-aware features by leveraging masks generated from the Segment Anything Model (SAM) and utilizing our novel contrastive learning objective based on hard pixel mining.
-The learned Gaussian features can be effectively clustered without further post-processing.
-This enables fast computation for further object-level editing, such as object removal, composition, and style transfer by manipulating the Gaussians in the scene.
-We further extend several dynamic novel-view datasets with segmentation benchmarks to enable testing of learned feature fields from unseen viewpoints.
-We evaluate SADG on proposed benchmarks and demonstrate the superior performance of our approach in segmenting objects within dynamic scenes along with its effectiveness for further downstream editing tasks.
-Our project page is available at: https://yunjinli.github.io/project-sadg/.
+Understanding dynamic 3D scenes is crucial for extended reality (XR) and autonomous driving.
+Incorporating semantic information into 3D reconstruction enables holistic scene representations, unlocking immersive and interactive applications.
+To this end, we introduce TRASE, a novel tracking-free 4D segmentation method for dynamic scene understanding.
+TRASE learns a 4D segmentation feature field in a weakly-supervised manner, leveraging a soft-mined contrastive learning objective guided by SAM masks.
+The resulting feature space is semantically coherent and well-separated, and final object-level segmentation is obtained via unsupervised clustering.
+This enables fast editing, such as object removal, composition, and style transfer, by directly manipulating the scene’s Gaussians.
+We evaluate TRASE on five dynamic benchmarks, demonstrating state-of-the-art segmentation performance from unseen viewpoints and its effectiveness across various interactive editing tasks.
+Our project page is available at: https://yunjinli.github.io/project-sadg/
 1 Introduction
-We live in a dynamic, three-dimensional world.
-Most of the time, we describe our environment at a high level of objects.
-As we move around and observe the scene from different viewpoints, we complete the geometry and perceive the colors contributing to our object representation.
-At the same time, we infer dynamic properties with object motion and deformations.
-Despite pose, color, and shape changes, the underlying semantic meaning stays constant.
-This intuition can serve as a strong constraint for neural 3D representations that have gained recent popularity and promote innovation in augmented reality, gaming, and autonomous driving interactive applications.
-In these technologies, seamlessly integrating semantic information into underlying representation ensures an immersive experience and gives users advanced control over the dynamic scene contents.
-Neural Radiance Fields (NeRFs) have revolutionized the field of 3D reconstruction and demonstrated excellent performance in synthesizing photorealistic novel views of a scene from any angle.
-The limitations of vanilla NeRFs promoted many follow-up works, including those that encompass semantic information by lifting 2D segmentation masks [24, 7, 19] or distilling features [21, 18] and those that faithfully represent dynamic scenes [41, 35, 27].
-Due to continuous implicit representation, these methods are computationally intensive as they require re-rendering from multiple views to ensure edit consistency, which hinders their applicability for interactive downstream tasks.
-Building on the strengths of NeRFs, Gaussian splatting (3DGS) [17] has introduced a powerful new approach that prioritizes efficiency and speed.
-Dynamic versions of 3DGS [30, 56] have offered faithful reconstruction of scenes with motion.
-These models can serve as a geometric prior for learning semantic features that are consistent in both space and time and open new possibilities for real-time interactive applications like scene editing and dynamic interactions.
-Towards this, a line of works has extended geometric properties of colored and explicit Gaussians with semantic information [6, 57], producing consistent novel view synthesis and segmentation.
-Nonetheless, these works are limited to static scenes and cannot handle objects in motion.
-There are only a few concurrent works [16, 25], which unify semantics and dynamics to comprehensively understand real-world environments.
-However, SA4D [16] relies on the supervision of object trackers and suffers from semantic identity conflicts in multi-view scenes.
-Moreover, their method is sensitive to mask noise and requires post-processing to remove degenerate Gaussians.
-At the same time, DGD [25] has a higher memory requirement and longer training time as they are learning 4D feature fields based on 512-dimensional CLIP [42] or 384-dimensional DINOv2 [34].
-Our work aims to tackle both single- and multi-view representations with fast rendering times, minimal post-processing, and low memory footprint.
-To this end, we introduce our novel framework, Segment Any Dynamic Gaussian Without Object Trackers (SADG), which effectively combines dynamic Gaussian splatting backbone [56] and semantic information without any reliance on trackers that maintain consistent object identities.
-To learn the spatio-temporal semantic field, we define a new contrastive objective based on hard positive and negative mining with the Segment Anything Masks (SAM) [20].
-We inherit the strengths of zero-shot 2D scene understanding of the foundation model to learn the semantically-aware feature field and produce consistent spatio-temporal segmentation.
-Our 32-dimensional features are compact and do not depend on features of a 2D segmentation model.
-Rendering time and storage requirements are almost unaffected, enabling real-time interaction and editing.
-The learned feature field can be effectively clustered without any further post-processing and utilized in a range of downstream editing tasks, including object style transfer [14], re-coloring, composition, and removal.
-Given an early stage of development and benchmarks for semantic segmentation for dynamic scenes, we propose to extend existing novel-view synthesis datasets with semantic benchmarks.
-The proposed evaluation protocol includes single- and multi-view scenes with challenging motions and diverse contents, allowing extensive testing of semantically-aware latent representations from unseen viewpoints.
+Humans experience the world in motion, where objects persist over time despite changes in pose, shape, and appearance.
+Our ability to recognize objects remains stable, even as we move and observe them from different vantage points.
+This phenomenon suggests the existence of neural 3D representations, which preserve semantic consistency across both space and time.
+Such representations are central for applications in augmented reality, gaming, and autonomous systems, where interactive manipulation of dynamic scenes requires object-level segmentation that is spatio-temporally consistent and efficient.
+Neural Radiance Fields (NeRFs) [31] have significantly advanced 3D reconstruction by enabling photorealistic novel view synthesis, but their implicit nature makes them computationally too expensive for real-time interaction.
+Gaussian Splatting (3DGS) [20] has recently emerged as a powerful alternative, achieving real-time rendering with explicit and compact scene representations.
+Dynamic extensions of 3DGS [28, 52] allow faithful reconstruction of moving scenes, but they lack built-in object segmentation and rely purely on geometry and color.
+While several works have integrated semantic information into static 3D Gaussian representations [6, 53], they fail to handle motion, deformation, and viewpoint variation in dynamic scenes.
+Only a handful of works have explored the unification of semantics and dynamics in 3DGS [19, 24].
+SA4D [19] incorporates segmentation into dynamic Gaussian splatting, using video object trackers to match masks across views similar to Gaussian Grouping [53].
+However, approaches using these trackers often suffer from identity switches, causing inconsistent segmentation in multi-view settings - particularly in the presence of occlusions, non-rigid deformations, or rapid motion, as shown in Fig.
+3.
+DGD [24] addresses temporal consistency by distilling semantically-aware features from foundation models like CLIP [45] and DINOv2 [32], enabling tracking.
+Yet, these features are high-dimensional, leading to significant computational overhead from both the large memory requirements and slow training.
+In this work, we present TRASE, a tracking-free segmentation framework for dynamic scene understanding.
+TRASE learns temporally consistent segmentation features in a contrastive setting using only 2D binary masks.
+Unlike prior contrastive segmentation methods [12, 13, 54], which explicitly enforce cluster formation during training and promote statistical consistency of features, TRASE applies contrastive learning directly at the pixel-pair level in a fine-grained rendered feature space, deferring object clustering to a final post-training step.
+This design enables TRASE to learn a segmentation field with both high temporal consistency and sharp, well-defined object boundaries.
+It also effectively mitigates common artifacts, such as floaters and out-of-FoV errors, that frequently degrade prior contrastive learning approaches.
+To the best of our knowledge, we are the first to introduce a benchmark for segmentation in novel views of dynamic scenes.
+While established datasets like DAVIS [35, 36, 3, 4] and VOS [47, 50, 51] address dynamic video segmentation, none evaluate segmentation quality in unseen viewpoints.
+Moreover, our benchmark is significantly larger than the evaluation protocol suggested in [24], as our test sequences comprise single- and multi-view scenes and captures real-world scenarios.
+Beyond achieving state-of-the-art segmentation accuracy, TRASE is tailored for real-time interactive scene editing.
+Our compact 32-dimensional feature space requires minimal post-processing and integrates seamlessly into object manipulation tasks such as style transfer, recoloring, composition, and removal.
 Our contributions can be summarized as follows:
-We propose Segment Any Dynamic Gaussian Without Object Trackers (SADG), a new approach that achieves multi-view consistent segmentation of dynamic scenes without any tracking supervision.
-We leverage information from 2D masks to learn semantically-aware latent representation for dynamic scenes using a novel contrastive learning objective.
-We extensively evaluate our method on single- and multi-view scenes across five dynamic novel-view benchmarks and demonstrate state-of-the-art segmentation performance.
-We demonstrate the general nature of our feature space and apply it to several downstream tasks, including object removal, style transfer, and scene composition.
-We offer an interactive framework that allows to edit the scene with a simple mouse click or a text prompt. Our tool does not depend on input IDs or 2D strokes and operates in real-time.
+We propose TRASE, a new approach for multi-view consistent segmentation of dynamic scenes without tracking labels.
+To achieve this, we design a novel contrastive learning objective based on differentiable feature rendering and contrastive learning, ensuring segmentation stability across time and views.
+We offer a broad benchmark for segmentation in novel views of dynamic scenes, which encompass a broad spectrum of camera setups and captured scenarios.
+We extensively evaluate our method on five dynamic datasets and achieve state-of-the-art segmentation performance in both single- and multi-view settings.
+We demonstrate the applicability of our feature space to several scene editing tasks, including object removal, style transfer, and scene composition.
 2 Related Works
-2.1 Semantic Feature Learning
+2.1 3D Segmentation
+In static scenes.
 With the emergence of various vision foundation models, such as Segment Anything Model (SAM) [22], DINO [5], and DINOv2 [32], researchers start to integrate the semantics from these foundation models into their reconstruction.
@@
 GARField [21] introduces hierarchical grouping, leveraging the physical scales of 2D masks to group objects of different sizes and represent scenes at different granularity scales.
-Gaussian Grouping [57] is a pioneering Gaussian-based approach that utilizes consistent object IDs generated by video tracker DEVA [8] to render object IDs into camera views.
+Gaussian Grouping [53] first proposes to utilize consistent object IDs generated by video tracker DEVA [8] to render object IDs into camera views.
 Concurrent work such as SAGA [6] utilizes the masks generated from SAM, combined with scales obtained from 3D data inspired by GARField, to segment objects of various sizes.
-Notably, these methods are designed for static scenes.
-To date, only two concurrent work, SA4D [16] and DGD [25], addresses the segmentation task in dynamic 3DGS.
-DGD extends the rasterization pipeline of 3DGS to be able to render semantic feature of each Gaussian.
+GAGA [29] proposes a mask group identity assignment technique using a 3D-aware memory bank to compensate for inaccurate semantic labels and use group IDs as pseudo-labels for identity encoding.
+Our model implicitly learns multi-view consistent semantic Gaussian features using our designed contrastive loss in a single training stage without identity assignments.
+In dynamic scenes.
+To the best of our knowledge, only two prior works, SA4D [19], and DGD [24], address the segmentation task in dynamic 3DGS.
+DGD extends the rasterization pipeline of 3DGS to be able to render the semantic features of each Gaussian.
 The rendered Gaussian feature map is then trained with the supervision of DINOv2 [32] or CLIP [37] to mimic the output from their feature extractors.
-The biggest drawback of such supervision is anticipated longer training times as the dimensions of the features generated by DINOv2 and CLIP are quite large (384 and 512).
-Similar to Gaussian Grouping and its derivatives [11, 31], SA4D relies on video trackers to associate masks across different views to produce consistent object IDs.
-These approaches face significant limitations, as the mask association often fails to offer consistent object IDs in dynamic scenes, such as the oven-mitts sequence from HyperNeRF [36] as shown in Fig.
-2.
+The drawbacks of such supervision are longer training times and bigger 3DGS models, as the dimensions of the features generated by DINOv2 and CLIP are quite large (384 and 512), and the Gaussian features are also stored.
+Following the identity encoding concept from Gaussian Grouping and its derivatives [10, 29], SA4D introduces a temporal identity feature field.
+As the method relies on consistent object IDs for supervision, it is sensitive to temporal tracking inaccuracies due to occlusion or fast motion.
+A concurrent work, Split4D [17], follows several proposed design choices, including single-frame supervision and a contrastive learning objective, and utilizes a recent 4D reconstruction model to achieve 4D segmentation.
+Our TRASE does not rely on tracking labels and effectively handles single- and multi-view data with fast rendering times, minimal post-processing, and low memory footprint.
+2.2 Contrastive Learning for 3D Segmentation
+A number of recent works define multi-view 3D segmentation as a contrastive learning objective.
+Contrastive Gaussian Clustering (CGC) [12] training aims to maximize intra-cluster similarity in the 2D rendered feature space by enforcing the feature’s proximity with its corresponding cluster’s mean feature and ensuring its spatial consistency.
+OmniSeg3D [54] follows a similar CGC objective and supervises 3D point clustering using SAM-based image patches.
+ContrastiveLift [13] proposes a slow-fast contrastive fusion strategy, where a slow network aggregates global object-level information while a fast network refines per-point features.
+OpenGaussian [46] enforces intra-cluster similarity and inter-cluster dissimilarity of rendered features and further trains their discretized version.
+Split4D [17] utilizes the same objective as [12] and additionally enforces the features to align with DINOv2 [32] in the early training stage.
+Our TRASE proposes a contrastive objective based on soft sample mining of pixel pairs, which focuses on only imperfect features and allows us to learn the segmentation field effectively in 4D.
+No costly distillation is required as we supervise the training with binary 2D masks.
 2.3 Scene Editing Applications
-Learning semantically-aware latent space opens opportunities for interactive downstream applications, allowing users to manipulate scenes on an object level.
-Several works proposed editing techniques for radiance field representations [13, 49], which are limited to a single object.
-On the other hand, scene-based methods [53, 24] learn a separate radiance field per object and render the full scene by combining learned opacities along the viewing rays.
-While the structure is modular, its editing application is limited to a pre-defined number of objects.
-Our method is guided by user interaction and is capable of operating on any object in the scene.
-Another line of work distills features such as DINO [34] into volumetric representation and shows editing capabilities on learned continuous feature field [48].
-Due to the implicit representation, scene editing requires a re-assessment of the final consistency by rendering from novel scenes, which impedes the model’s applicability to real-time interactive applications.
-At the same time, Gaussian splatting methods mitigate the runtime limitation of NeRFs with their explicit point-based representation.
-SAGA [57] proposes Local Gaussian Editing scheme, which localize the Gaussians and manipulate them according to the editing task.
-Feature3DGS [60] distills features from 2D foundation models such as SAM [43] and LSeg [26] enabling promptable and language-driven 3D edits.
-Nonetheless, these methods operate only on static scenes, while our method is capable of editing dynamic objects and ensuring temporal consistency of the edits.
-DGD [25] shows semantic texture editing performance by utilizing a pre-trained diffusion model and fine-tuning dynamic 3DGS representation with the reconstruction loss.
-On the other hand, we demonstrate a variety of editing tasks, including style transfer, object removal, and composition.
-Despite offering several scene editing opportunities, SA4D [16] has limited user interactivity and requires prior knowledge of object identities.
-Our framework contains a user-friendly graphical interface that allows operation with a simple text prompt or a mouse click.
+Learning a semantically aware latent space enables interactive scene editing at the object level.
+Existing radiance-field editing methods are either restricted to single objects [16, 45] or rely on per-object scene decompositions [49, 23], which limit editing to a fixed set of instances.
+Other approaches, such as DINO [32], distill features into volumetric fields [44], but their implicit representations require expensive re-rendering for consistency, hindering real-time interaction.
+Gaussian-based methods [20] alleviate NeRF runtime issues via explicit point-based primitives: SAGA [53] enables local Gaussian editing, and Feature3DGS [56] leverages 2D foundation models like SAM [38] and LSeg [25] for promptable, language-guided 3D edits, yet all focus on static scenes.
+In contrast, our method edits dynamic objects with temporally consistent results and supports diverse tasks, including style transfer, object removal, and composition [24].
+While SA4D [19] offers semantic scene editing, it requires prior object labels and limited interaction.
+In contrast, our framework provides a user-friendly graphical interface that supports simple text prompts and mouse-based selection.
 3 Method
-In this section, we introduce our novel framework SADG.
+In this section, we introduce our novel TRASE.
 As illustrated in Fig.
-1 the pipeline comprises two main components: dynamic geometry reconstruction (Sec.
+2, the pipeline comprises two main components: dynamic geometry reconstruction (Sec.
 3.1) and Gaussian feature learning (Sec.
@@
 We follow the approach proposed in [52] and describe it briefly for completeness.
-Unlike Dynamic3DGS [30] model, where the reconstruction is stored per frame, resulting in high memory consumption, an MLP learns per-Gaussian deformation with respect to the static canonical space , which is defined as
+Unlike Dynamic3DGS [28] model, where the reconstruction is stored per frame, resulting in high memory consumption, an MLP learns per-Gaussian deformation with respect to the static canonical space , which is defined as for .
 is the center position of the i-th Gaussian, is the quaternion representing the rotation, is the scaling vector, refers to the opacity, and is the Spherical Harmonic Coefficients encoding the color information () as per [52].
@@
 3.2 Gaussian Feature Learning
-Problem statement Most of the existing works [16, 57] utilize a video tracker to provide consistent object mask IDs across views and render the segmentation field directly by predicting the per-pixel probability of different object classes.
-However, such approaches have a significant bottleneck, namely their reliance on perfect object mask association across views.
-When mask inconsistencies of the same object occur between views, the entire optimization pipeline can break down, leading to sub-optimal results or failure.
-Therefore, we propose a method that aims to learn the 3D semantics of objects and leverage them directly for downstream tasks across different views.
-In other words, instead of predicting per-pixel probabilities during rendering, we focus on integrating semantic information into Gaussians and leverage semantic constancy for learning multi-view consistent spatio-temporal feature field.
-Representation Given 4D reconstruction, we incorporate semantic information into the scene.
-Specifically, we extend its -th Gaussian with a 32-dimensional semantic feature vector , the so-called “Gaussian feature”.
-The semantically-aware feature field of a scene represented with Gaussians can be expressed as
-Rendering Gaussian Features The learned Gaussian features can be rendered similarly to [17, 60] with point-based -rendering [23, 22].
-Given a sorted subset of Gaussians that are related to pixel , the resulting rendered features can be computed as follows:
-Existing methods [6] use smooth Gaussian features instead, as they help to filter out noisy 3D Gaussian features with high similarity scores corresponding to unrelated objects.
-The smooth Gaussian feature for the -th Gaussian is computed as per Eq.
-6, where denotes a set of nearest neighbors [9] of the -th Gaussian based on its 3D position.
-We adopt this convention in our work and utilize smooth feature vectors in the rendering process (Eq.
-5).
-Our contrastive semantically-aware learning objective
-The generated set of SAM masks comprise binary masks of a given image , where each mask corresponds to a different object.
-In practice, we do not consider all masks from SAM to alleviate memory and computational requirements.
-Instead, we randomly sample masks for further usage.
-We define the pixel-mask correspondence vector in Eq.
-7, which captures mask information for a certain pixel coordinate (, ) across all binary masks in .
+Summary.
+Our feature learning augments each 3D Gaussian with a 32-dimensional feature vector, enabling segmentation without explicit tracking labels.
+It is worth noting that, in contrast to view-dependent spherical encoding, our semantic feature vector is consistent across different views and time.
+Thanks to differentiable feature rendering [56] and SAM [22] masks, we construct and explicitly enforce pixel-mask associations in the Gaussian feature field.
+Our novel contrastive learning objective is designed to be efficient by sub-sampling and effective by integrating SAM [22] masks.
+Using only 2D guidance and without any object labels, Gaussians are trained to separate per-pixel rendered features from different segments and align those within the same segment.
+To achieve this, we design a soft-mining objective.
+Unlike hard mining, which selects pairs strictly by similarity, soft mining uses a relaxed, existential rule that increases pixel coverage and makes training less sensitive to noisy feature similarities.
+Finally, the learned features are globally clustered, producing multi-view consistent segmentation.
+Our notation.
+A generated set of SAM masks comprise binary masks of a given image , where each mask corresponds to a different segment.
+With this, we define the pixel-mask correspondence vector , which captures information across all binary masks in for a certain pixel coordinate (, ).
 This vector is similar to a one-hot encoding technique; however, there is no guarantee that a pixel belongs to a single mask due to 2D segmentation inaccuracies.
@@
 4.
-where and .
-In practice, we only sample pixels from an image such that . Mathematically, we define the sampled pixel coordinates as a set .
-The pairwise similarities of the sampled pixel-mask correspondence vectors can be retrieved by computing the Gram matrix of their stacked matrix .
+The similarities of pixel-mask vectors for a pair of pixels (, ) can be retrieved by computing the Gram matrix of their stacked matrix , where refers to the number of sampled pixels.
 All entries of the resulting Gram matrix are then bounded to either 0 (negative pair) or 1 (positive pair) to form a mask-based correspondence matrix among sampled pixels, formally defined as
-Additionally, we formulate a correspondence matrix for the sampled rendered features .
-Analogous to the mask-based correspondence matrix , we compute a feature-based Gram matrix, capturing the pairwise similarities between the features as follows:
-where is defined as follows:
-Finally, we have all the necessary ingredients to formulate the contrastive learning objective based on a hard positive and negative mining scheme.
-Guided by the mask-based correspondence matrix, we enforce features to be close for pixels from the same object mask and far apart for the distinct objects.
-Furthermore, we concentrate on hard pixel pairs.
-Specifically, we sample positive samples with low similarity scores and negative samples with high similarity scores.
-With this, we formally define the optimization objective as
-Note that is an indicator function.
-It returns 1 when , and 0 otherwise.
-stands for hard positive which is defined as and refers to a hard negative which is defined as .
-Here, and are the positive and negative thresholds for pairwise similarity respectively.
-As both and are symmetric matrices following the property of the Gram matrix, we only consider entries in the upper triangular part and exclude the diagonal to avoid repeated computation.
-We show rendered Gaussian features learned with our contrastive scheme in Fig.
-1.
-For demonstration, they are compressed into via PCA [33], allowing each rendered feature to be represented with RGB color.
-Gaussian Feature Clustering To enable interactive applications with the learned 4D representation, we deploy the DBSCAN algorithm [12] to cluster the Gaussian features into several groups and generate the so-called Gaussian Clusters as defined in Eq.
-12.
-provides the corresponding cluster ID for -th Gaussian feature:
-Note that refers to the total number of clusters. We demonstrate them visually in Fig. 1. As can be observed, they represent distinct objects in 4D space.
-Since a typical scene contains a large number of Gaussians (more than 300k), we adopt the sub-sampling scheme proposed by [6], where only 2% of the Gaussian features are used for clustering.
-We propagate the cluster IDs to the remaining of the Gaussians by computing the feature similarity between cluster-representative Gaussians.
-In the supplementary material, we demonstrate that the performance is not affected by a significant reduction in the number of Gaussians contributing to the generation of clusters.
+Analogous to the binary mask-based Gram matrix , we compute a feature-based matrix , capturing the pairwise similarities between the rendered features .
+Following common practices [6], we consider a smooth version of Gaussian features, as it is more robust than considering individual features, which might have high similarity scores corresponding to unrelated objects.
+The smooth Gaussian feature is computed as the mean feature of its nearest neighbors [9] based on the Gaussian’s 3D position.
+Our learning objective.
+To robustly learn multi-view coherent feature embeddings, we adopt a soft-mined contrastive learning objective.
+Unlike standard hard mining, which selects positive or negative pairs strictly based on their individual similarity values, soft mining employs an existential criterion: a positive pixel pair (,) is considered into final positive loss computation if there exists any positive pair involving whose similarity falls below the positive threshold .
+Similarly, a pair is included in the negative loss if there exists any negative pair involving whose similarity exceeds the negative threshold .
+By design, the soft-mined strategy broadens pixel coverage during training, in contrast to naive hard sampling that typically concentrates on a sparse set of imperfect pixel pairs and leaves many pairs untrained.
+We further restrict the selection to non-redundant pairs by considering only the upper-triangular portion of the matrices ().
+Formally, the soft-mined masks are defined as
+The positive and negative losses are then computed as weighted sums over these masks:
+where is a normalization factor (e.g., the total number of pixel pairs), is a pixel-weight matrix similar to [6], and ensures that only positive contributions are considered for negative pairs.
+The final loss is then defined in Eq.
+8 with an extra regularization term to stabilize the feature learning.
+Inference.
+To enable interactive applications with the learned 4D representation, we deploy DBSCAN algorithm [11] to cluster the Gaussian features into several groups and generate the so-called Gaussian clusters.
+In contrast to the K-Means algorithm [27, 30], it does not require any prior knowledge of the number of clusters in a scene.
+As can be seen in Fig.
+2, clusters represent distinct objects, which verifies accurate alignment of our learned feature space with the object-level semantic field.
+We further observe that the object boundaries can be effectively refined by applying a simple similarity threshold, which prunes Gaussians with features that are less similar to the corresponding cluster’s mean.
+This simple technique enhances the method’s performance in challenging segmentation scenarios, such as those involving occlusions and affected by sparse observations.
 4 Experimental Results
-We evaluate the performance of SADG across several well-known datasets to validate its effectiveness.
-We begin by introducing the datasets in Sec.
-4.1.
-Since there are no established benchmarks for validating the segmentation performance on novel views in dynamic scenes, we provide details on how we extend these different datasets with segmentation benchmarks (Sec.
-4.2).
-Next, we explore the implementation details and various configurations used to train and test methods on these benchmarks (Sec.
-4.3).
+We first provide details of our proposed 4D segmentation benchmark in Sec.
+4.1 and outline the evaluation protocol in Sec.
+4.2 .
 We present segmentation results of our model, along with both quantitative and qualitative comparisons to other related works [6, 53, 24] in Sec.
-4.5.
-Last but not least, we demonstrate capabilities of our learned semantic representation on a number of downstream tasks including scene editing and selection in our a graphical user interface (Sec.
-4.6).
-4.1 Datasets
+4.3.
+Last but not least, we demonstrate the capabilities of our learned semantic representation on a number of downstream tasks, including scene editing and selection via our designed graphical user interface (Sec.
+6).
+Please refer to the supplementary material for implementation details.
+4.1 4D Segmentation Benchmark
+We manually annotate dynamic sequences of existing novel-view synthesis datasets and present their segmentation-aware versions: NeRF-DS-Mask, HyperNeRF-Mask, Neu3D-Mask, Immersive-Mask, and Technicolor-Mask.
+Utilizing the powerful SAM2 model [38], which provides consistent object masks across video sequences, we define the objects of interest for each sequence to evaluate their segmentation performance in our model and manually refine them.
+Qualitative examples of the augmented sequences, together with a description of the raw datasets, can be found in the supplementary material.
 | NeRF-DS-Mask | HyperNeRF-Mask | Neu3D-Mask | Immersive-Mask | Technicolor-Mask | ||||||
 | Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
-Gaussian Grouping [57] | 0.8351 | 0.9729 | 0.8341 | 0.9771 | 0.8864* | 0.9932* | 0.7673* | 0.9776* | 0.7979* | 0.9800* |
+| GG [53] | 0.8351 | 0.9729 | 0.8341 | 0.9771 | 0.8864* | 0.9932* | 0.7673* | 0.9776* | 0.7979* | 0.9800* |
 | SAGA [6] | 0.8072 | 0.9644 | 0.7660 | 0.9579 | 0.6941 | 0.9516 | 0.7395 | 0.9747 | 0.7913 | 0.9792 |
 | DGD [24] | 0.7125 | 0.9551 | 0.8297 | 0.9777 | 0.7721 | 0.9851 | 0.7981 | 0.9850 | 0.7791 | 0.9728 |
-SADG (Ours) | 0.8719 | 0.9826 | 0.8663 | 0.9845 | 0.9022 | 0.9945 | 0.9234 | 0.9945 | 0.9308 | 0.9917 |
-NeRF-DS dataset [52] contains scenes with challenging transparent and shiny moving objects recorded by two monocular cameras.
-Five sequences are used for reconstruction and evaluation including as_novel_view, basin_novel_view, cup_novel_view, press_novel_view, and plate_novel_view.
-HyperNeRF dataset [36] is recorded using a single monocular camera .
-The camera moves while capturing the dynamic scenes involving different human-object interactions.
-In our experiments, we use ten sequences: americano, split-cookie, oven-mitts, espresso, chickchicken, hand1-dense-v2, torchocolate, slice-banana, keyboard, and cut-lemon1.
-Neu3D dataset [27] contains videos captured by a multi-view rig with to cameras each at 30 FPS.
-The scenes involve a person performing cooking tasks in the kitchen.
-In total, five scenes are used in our experiments: coffee_martini, cook_spinach, cut_roasted_beef, flame_steak, and sear_steak.
-Google Immersive dataset [2] consists of videos captured by a 46-camera rig at 30 FPS.
-The cameras are fish-eye cameras and are distributed on the surface of a hemispherical, 92cm diameter dome.
-Four scenes are used in our experiments: 01_Welder, 02_Flames, 10_Alexa_1, and 11_Alexa_2.
-Technicolor Light Field dataset [44] captures scenes using a camera array at 30 FPS.
-We use the undistorted images provided by the authors, and four scenes are used in our experiments: Birthday, Fabien, Painter, and Theater.
-4.2 Segmentation Benchmarks
-To date, while numerous benchmarks exist for dynamic video segmentation, such as DAVIS [38, 40, 3, 4] and VOS [51, 54, 55], there are still no well-established segmentation benchmarks for dynamic novel view synthesis.
-Such benchmarks are crucial for evaluating a comprehensive understanding of scenes, both geometrically and semantically.
-Existing datasets, such as Replica [47], ScanNet [10], and ScanNet++ [58], include ground-truth labels for evaluating panoptic segmentation performance, but they do not cover dynamic sequences.
-To address this gap, we manually annotate dynamic sequences in our experiments and propose the NeRF-DS-Mask, HyperNeRF-Mask, Neu3D-Mask, Immersive-Mask, and Technicolor-Mask benchmarks.
-Utilizing the powerful SAM2 model [43], which provides consistent object masks across video sequences, we manually define the objects of interest for each sequence to evaluate their segmentation performance in our model.
-Examples of these mask objects can be seen in the supplementary.
-4.3 Implementation Details
-We employ SAM [20] to generate the anything-masks for Gaussian feature learning.
-However, the resulting masks are memory-inefficient because they are stored in a dense format () with a lot of pixels with zero values.
-To optimize memory usage, we save the masks as a bit array.
-Given a permuted tensor version of an RGB image , SAM generates masks .
-We first flatten , then convert flattened ()-dimensional array into a bit-array.
-The final output is stored in a dictionary with the required information to re-create the original mask tensor: a number of masks , image dimensions , and the compressed bit array, i.e., .
-This approach significantly reduces the storage used and is particularly useful for multi-view datasets, where numerous masks from different views are generated.
-For instance, a raw mask generated from an image in coffee_martini occupies around 80 MB, but the dictionary representation is significantly reduced to approximately 1 MB.
-Due to space limits, we provide training details for different datasets in the supplementary.
+| SA4D [19] | 0.6740 | 0.9360 | 0.7371 | 0.9616 | 0.8832* | 0.9391* | 0.7987* | 0.9835* | 0.8271* | 0.9734* |
+| CGC [42] | 0.8100 | 0.9718 | 0.8157 | 0.9768 | 0.8754 | 0.9927 | 0.8856 | 0.9925 | 0.8806 | 0.9812 |
+| OpenGaussian [46] | 0.6939 | 0.9564 | 0.6311 | 0.9411 | 0.8178 | 0.9899 | 0.7120 | 0.9793 | 0.8390 | 0.9694 |
+| TRASE (Ours) | 0.8768 | 0.9831 | 0.8663 | 0.9845 | 0.9022 | 0.9945 | 0.9234 | 0.9945 | 0.9308 | 0.9917 |
 4.2 Evaluation Protocol and Baselines
-Given ground truth masks we use Mean Intersection over Union (mIoU) and Mean Pixel Accuracy (mAcc) to quantitatively evaluate segmentation performance of the models.
-We compare our model with SAGA [6], Gaussian Grouping [57] and DGD [25].
-Note that both SAGA and Gaussian Grouping are designed for static scenes and do not include a module to encode temporal information.
-To address this limitation, we integrate our 4D reconstruction backbone into their pipeline, extending their approach to effectively handle dynamic scenes.
-Moreover, Gaussian Grouping needs the precomputed object IDs from DEVA [8], whose labels are inconsistent for multi-view camera setup.
-Thus, we train their semantic module on a monocular stream from the camera closest to the test view.
-Furthermore, all reported baselines utilize the same pre-trained 4D reconstruction, with semantics learned according to each pipeline.
-This allows us to isolate and evaluate the performance of each model based on its semantic learning capabilities.
-After training, we manually perform click prompts in the rendered novel views of each scene, which are trained by different approaches to select the objects of interest defined in the benchmark.
+We compare our model with SAGA [6], Gaussian Grouping (GG) [53], SA4D [19], and DGD [24].
+Additionally, we add evaluation with the contrastive learning methods such as Contrastive Gaussian Clustering (CGC) [12] and OpenGaussian [46].
+Note that SAGA, Gaussian Grouping, CGC, and OpenGaussian are designed for static scenes and do not include a module to encode temporal information.
+To evaluate them on the 4D segmentation task, we integrate our 4D reconstruction backbone into their pipeline, extending their approach to effectively handle dynamic scenes.
+GG and SA4D require video object tracking labels for their supervision.
+To enable evaluation of the multi-view sequences from our benchmarks, we train the models on the camera closest to the test view (cam5 for Neu3D, cam2 for Immersive, and cam1 for Technicolor).
+After training, we manually perform click prompts in the rendered novel views of each scene trained by different approaches to select the objects of interest defined in the benchmark.
 Apart from the quantitative results, we also present qualitative results from various benchmarks, as certain insights cannot be fully captured through numerical evaluations alone.
-Since the official source code for SA4D [16] is not available, we report the comparison based on unofficial implementation in the supplementary.
+Following common practices, we use Mean Intersection over Union (mIoU) and Mean Pixel Accuracy (mAcc) to quantitatively evaluate the performance of the models.
 4.3 Segmentation Results
+| Preproc | Training | Total (min) | Storage (MB) | |
+| DGD [24] | 0.08 | 103 | 103 | 285 |
+| TRASE (Ours) | 16 | 16 | 32 | 60 |
 As shown in Tab.
-1, SADG consistently outperforms all other methods across all benchmarks in terms of average mIoU and mAcc.
-We also select some sequences to visually illustrate the quality of the segmentation of each model in Fig.
+1, TRASE achieves state-of-the-art performance across all benchmarks, surpassing other segmentation methods.
+Please refer to the supplementary material for details on each sequence of the benchmark.
+Further, we select some sequences to illustrate the quality of the segmentation of each model in Fig.
 5.
-These comparisons reveal that SADG qualitatively surpasses existing models and exhibits multi-view consistent segmentation performance.
-Notably, SADG maintains strong generalization to test camera views in multi-view sequences such as flame_steak, Painter, and 11_Alexa_2.
+These comparisons reveal that TRASE qualitatively surpasses existing models and exhibits multi-view consistent segmentation performance.
+Notably, TRASE maintains strong segmentation consistency to test camera views in multi-view sequences such as flame_steak, Painter, and 11_Alexa_2.
 This robust performance can be attributed to the compact Gaussian features learned through the proposed Gaussian feature training.
@@
 However, determining a suitable threshold that generalizes across various scenes is neither straightforward nor intuitive.
-SADG employs clustering based on the Gaussian features to group the Gaussians into several segments.
-As this operation is done in 3D space, SADG achieves cross-view consistency and does not need to associate objects from different views.
+TRASE employs clustering based on the Gaussian features to group the Gaussians into several segments.
+As this operation is done in 3D space, TRASE achieves cross-view consistency and doesn’t need to associate objects from different views.
 This advantage is evident in cut-lemon1 example in Fig.
-5, where wrongly associated object IDs can cause segmentation failure for Gaussian Grouping.
-4.6 Scene Editing Applications
+5, where wrongly associated object IDs can cause segmentation failure for [53].
+5 Ablation Studies
+5.1 Time and Storage Analysis
+We perform a comprehensive time and storage analysis of TRASE and DGD [24] using an RTX 3080 and an Intel i7-12700 on the NeRF-DS [48] dataset.
+The results are shown in Tab.
+2, where “Pre-Proc” refers to the generation of DINOv2 384-dim features for DGD and SAM masks for TRASE.
+Rendering 384-dim features into 2D is computationally inefficient, significantly slowing down DGD’s training process.
+TRASE uses 32-dim Gaussian features, which can be rendered much more efficiently.
+As a result, TRASE achieves over 3 faster total time than DGD while only requiring 4.5 less space in storage.
+5.2 Soft Sample Mining
+We visually demonstrate the effectiveness of our soft-mined pixel-pair sampling strategy in comparison to other contrastive objectives: sampling all positive–negative pairs (all) and sampling only pairs that individually satisfy the thresholds (hard).
+We include contrastive-based baselines such as CGC [12] and OpenGaussian [46] for completeness.
+By investigating the out-of-FoV regions, i.e., 3D regions outside the observable space of train and test views, we observed floaters in the segmentation by our method without soft selection and other contrastive baselines, as shown in Fig.
+6.
+Our mining approach effectively eliminates these floaters, leading to more stable and reliable segmentation.
+It is worth noting that no post-processing filtering is applied to ensure a fair comparison.
+5.3 Single-frame Mask Supervision
+Inspired by classical tracking methods [15] and recent work linking 3DGS dynamics to pixel motion [14], we investigate the impact of incorporating temporal constraints on our single-frame mask supervision.
+Specifically, we compare our standard single-frame approach with a temporal extension, where masks are propagated across consecutive frames to encourage segmentation consistency over time.
+To implement this, we estimate optical flow between consecutive training timestamps and using differentiable rendering with accumulated 2D Gaussian displacements [14], and warp mask pixels from to .
+Quantitative results are presented in Tab.
+3.
+As observed, the multi-frame baseline does not provide significant improvements over the single-frame supervision, despite the doubled training cost, as an additional 2D flow and all associated parameters are rendered.
+This validates our design choice to rely on single-frame mask supervision, which achieves strong segmentation performance efficiently.
+| HyperNeRF-Mask | Immersive-Mask | Technicolor-Mask | ||||
+| Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
+| TRASE (multi-frame) | 0.8636 | 0.9833 | 0.8968 | 0.9933 | 0.9309 | 0.9913 |
+| TRASE (Ours) | 0.8663 | 0.9845 | 0.9234 | 0.9945 | 0.9308 | 0.9917 |
+| NeRF-DS-Mask | HyperNeRF-Mask | Neu3D-Mask | |||||
+| Method | Filtering | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
+| SAGA [6] | ✘ | 0.8072 | 0.9644 | 0.7660 | 0.9579 | 0.6941 | 0.9516 |
+| ✓ | 0.8421 | 0.9778 | 0.8223 | 0.9779 | 0.8026 | 0.9853 | |
+| TRASE | ✘ | 0.8760 | 0.9829 | 0.8368 | 0.9790 | 0.8738 | 0.9923 |
+| TRASE (Ours) | ✓ | 0.8768 | 0.9831 | 0.8663 | 0.9845 | 0.9022 | 0.9945 |
+5.4 Post-processing Filtering
+We provide an ablation study of our method with respect to the filtering step, which removes Gaussians whose features are too dissimilar to the object’s mean.
+As shown in Tab.
+4, the proposed filtering step enhances the segmentation quality of our method.
+Particularly, artifacts on the object’s boundaries are effectively removed, mitigating the inaccuracies in input masks or the effects of occlusions.
+We further demonstrate the advantage of our learned feature field by showing that while the post-processing step enhances segmentation accuracy, it is not the sole factor driving our superior performance.
+We also note that this technique is applicable only to certain baselines such as SAGA [6].
+For example, methods like CGC [12] and DGD [24] already incorporate a similarity threshold during inference.
+OpenGaussian [46] gets the assigned class ID for each Gaussian from their codebook.
+By contrast, SA4D [19] and Gaussian Grouping [53] learn an identity field, making our post-processing inapplicable to them.
+In Tab.
+4 and Fig.
+7, we quantify and illustrate the effectiveness of applying our proposed filtering step to SAGA [6].
+However, because the features learned under the baseline’s hard-contrastive objective remain suboptimal, the segmentation quality still exhibits more artifacts compared to TRASE.
+6 Editing Applications
 In addition to accurate segmentation capabilities and multi-view consistent rendered masks, we demonstrate the effectiveness of our learned semantic feature field in a range of downstream tasks.
 Specifically, we focus on interactive editing applications and develop a graphical user interface (GUI) to grant full user control over the representation.
-Thanks its explicit nature and our semantic abstraction to Gaussian clusters, we achieve real-time operation capability in handling many editing tasks.
+Thanks to its explicit nature and our semantic abstraction to Gaussian clusters, we achieve real-time operation capability in handling many editing tasks.
 We provide snapshots of the GUI in the supplementary.
-Besides simple click prompts in the novel view of dynamic scenes for object selection, we also support text-based prompts to make object selection more intuitive.
+In addition to simple click prompts, we also support text-based prompts to make object selection even more intuitive.
 To enhance 3D content editing in dynamic scenes, we integrate style transfer capabilities for the selected objects.
 By segmenting objects directly in 3D, we facilitate scene composition and object removal through straightforward manipulation of the selected Gaussians.
-Due to space limit, we shall provide more qualitative results in the supplementary.
-In the following we describe the implementation details of editing tasks based on our semantically-aware feature field.
+Due to the space limit, we shall provide more qualitative results in the supplementary.
+In the following, we describe the implementation details of editing tasks based on our semantically-aware feature field.
 Text Prompt.
-Inspired by Gaussian Grouping [57], we employ Grounding DINO [28] to generate 2D mask based on the text prompt in the novel view.
+Inspired by Gaussian Grouping [53], we employ Grounded SAM [39] to generate a 2D mask based on the text prompt in the novel view.
 The 2D mask is reprojected into the 3D space with the rendered depth information.
 Finally, we associate each reprojected point with its nearest Gaussian in the 3D scene and retrieve its corresponding cluster.
-If the number of associated Gaussian of a given cluster is more than a threshold, the cluster would be selected.
+If the number of associated Gaussians of a given cluster is more than a threshold, the cluster would be selected.
 The example of text prompts on different sequences is shown in Fig.
@@
 The scales and coordinate transformation need to be adapted.
-We show some examples on scene composition in Fig.
-6.
+We show some examples of scene composition in Fig.
+8.
 Object Style Transfer.
@@
 Object Removal.
-We can also remove object by simply deleting Gaussians belonging to the specific cluster as illustrated in Fig.
+We can also remove objects by simply deleting Gaussians belonging to the specific cluster as illustrated in Fig.
 8 (d).
-It is worth noting that no inpainting is required due to the underlying 3D representation, where exposed parts are learned by multi-view observations.
+It is worth noting that an additional inpainting step may be required, where the newly exposed parts have not been learned from multi-view observations.
 7 Conclusion
 We introduced a novel framework for dynamic scene understanding, which enables multi-view consistent segmentation without any object tracking supervision.
-Our SADG effectively combines dynamic 3D Gaussian Splatting 3DGS [56] and 2D SAM [20] masks in a contrastive learning objective, which lifts semantic information into 3D space and learns expressive Gaussian features based on hard pixel cases.
+Our TRASE effectively combines dynamic 3D Gaussian Splatting 3DGS [52] and 2D SAM [22] masks in a contrastive learning objective, which lifts semantic information into 3D space and learns expressive Gaussian features based on soft sample mining.
 This results in cross-view consistency when rendering segmented objects, enhancing the quality and coherence of object segmentation across different views.
-Evaluated on various novel-view datasets, SADG shows superior performance both quantitatively and qualitatively.
-We further demonstrated the effectiveness of the learned feature field on downstream editing tasks such as point and text prompts, style transfer, object removal, and scene composition.
+Evaluated on various novel-view datasets, TRASE shows superior performance both quantitatively and qualitatively.
+We further demonstrate the effectiveness of the learned feature field on downstream editing tasks such as point and text prompts, style transfer, object removal, and scene composition.
 Our approach sets a strong foundation for further research into dynamic scene understanding and scene editing, especially in complex multiview scenarios.
+We intend to release our code and benchmarks for future developments.
 Supplementary Material
-In the supplementary, we provide further implementation details in Sec.
+In the supplementary, we provide several examples from our segmentation benchmark, along with dataset details (Sec.
+8).
+Further implementation details are described in Sec.
 9 and Sec.
-7.
+10.
 A range of ablation studies on different components of our pipeline is given in Sec.
-8.
-We compare our SADG with the concurrent work on 4D segmentation and editing, Segment Any 4D Gaussians (SA4D), in Sec.
+11.
+In Sec.
+12 we address several limitations of our method and potential directions for future work.
+We provide additional visualizations of our learned feature space in comparison to other contrastive-based segmentation methods in Sec.
+13 and Sec.
+14.
+In Sec.
+15 we show more examples on how our soft mining successfully mitigates floaters.
+Further, we present detailed quantitative results on all benchmarks in Sec.
+17.
+We finally demonstrate qualitative results of our method on segmentation (Sec.
+18, Sec.
+19) and editing (Sec.
+16).
+8 Segmentation Benchmark
+As discussed in Sec.
+4.1 (main), we employ SAM2 [38] to manually select objects of interest to form our segmentation benchmarks for dynamic scene novel view synthesis.
+The examples are illustrated in Fig.
 9.
-Several examples from our segmentation benchmark are given in Sec.
-11.
-We present detailed quantitative results on all benchmarks in Sec.
-10.
-We finally provide qualitative results of our method on segmentation (Sec.
-12, Sec.
-14) and editing (Sec.
-13).
+We describe the details of the mask storage for training and datasets adapted for our segmentation benchmark in the following.
+Mask Storage.
+We employ SAM [22] to generate the anything-masks for Gaussian feature learning.
+However, the resulting masks are memory-inefficient because they are stored in a dense format () with a lot of zero-pixels.
+To optimize memory usage, we save the masks as a bit array.
+Given a permuted tensor version of an RGB image , SAM generates masks .
+We first flatten , then convert flattened ()-dimensional array into a bit-array.
+The final output is stored in a dictionary with the required information to re-create the original mask tensor: a number of masks , image dimensions , and the compressed bit array, i.e., .
+This approach significantly reduces the storage used and is particularly useful for multi-view datasets, where numerous masks from different views are generated.
+For instance, a raw mask generated from an image in coffee_martini occupies around 80 MB, but the dictionary representation reduces to about 1 MB.
+NeRF-DS.
+[48] contains scenes with challenging transparent and shiny moving objects recorded by two monocular cameras.
+Five sequences are used for reconstruction and evaluation including as_novel_view, basin_novel_view, cup_novel_view, press_novel_view, and plate_novel_view.
+HyperNeRF. [33] is recorded using a single monocular camera
+.
+The camera moves while capturing the dynamic scenes involving different human-object interactions.
+In our experiments, we use ten sequences: americano, split-cookie, oven-mitts, espresso, chickchicken, hand1-dense-v2, torchocolate, slice-banana, keyboard, and cut-lemon1.
+Neu3D.
+[26] contains videos captured by a multi-view rig with to cameras each at 30 FPS.
+The scenes involve a person performing cooking tasks in the kitchen.
+In total, five scenes are used in our experiments: coffee_martini, cook_spinach, cut_roasted_beef, flame_steak, and sear_steak.
+Google Immersive.
+[2] consists of videos captured by a 46-camera rig at 30 FPS.
+The cameras are fish-eye cameras and are distributed on the surface of a hemispherical, 92cm diameter dome.
+Four scenes are used in our experiments: 01_Welder, 02_Flames, 10_Alexa_1, and 11_Alexa_2.
+Technicolor Light Field.
+[40] captures scenes using a camera array at 30 FPS.
+We use the undistorted images provided by the authors, and four scenes are used in our experiments: Birthday, Fabien, Painter, and Theater.
 9 Training Implementation Details
@@
 We reuse the codebase from Deformable-3DGS [52] and extend the differential Gaussian rasterization pipeline to render the Gaussian features similar to [53].
-In the following, we introduce the detailed training setup for different datasets.
+In the following sections, we present the detailed training and evaluation setup for various datasets.
 9.1 NeRF-DS and HyperNeRF
@@
 | Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
-SADG (w/o filtering) | 0.8719 | 0.9826 | 0.8368 | 0.9790 | 0.8738 | 0.9923 | 0.9154 | 0.9942 | 0.9258 | 0.9904 |
-Ours (w/ filtering) | 0.8719 | 0.9826 | 0.8663 | 0.9845 | 0.9022 | 0.9945 | 0.9234 | 0.9945 | 0.9308 | 0.9917 |
-NeRF-DS-Mask | HyperNeRF-Mask | Neu3D-Mask | Immersive-Mask | Technicolor-Mask | ||||||
-Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
-SADG (K=0) | 0.8228 | 0.9706 | 0.8360 | 0.9787 | 0.8661 | 0.9920 | 0.8725 | 0.9909 | 0.9259 | 0.9908 |
-Ours (K=16) | 0.8719 | 0.9826 | 0.8368 | 0.9790 | 0.8738 | 0.9923 | 0.9154 | 0.9942 | 0.9258 | 0.9904 |
-SADG (K=32) | 0.8754 | 0.9828 | 0.8331 | 0.9780 | 0.8702 | 0.9922 | 0.9217 | 0.9944 | 0.9255 | 0.9904 |
+| TRASE (K=0) | 0.8228 | 0.9706 | 0.8360 | 0.9787 | 0.8661 | 0.9920 | 0.8725 | 0.9909 | 0.9259 | 0.9908 |
+| Ours (K=16) | 0.8760 | 0.9829 | 0.8368 | 0.9790 | 0.8738 | 0.9923 | 0.9154 | 0.9942 | 0.9258 | 0.9904 |
+| TRASE (K=32) | 0.8754 | 0.9828 | 0.8331 | 0.9780 | 0.8702 | 0.9922 | 0.9217 | 0.9944 | 0.9255 | 0.9904 |
 Dynamic Geometry Reconstruction.
@@
 For each image, we randomly sample 25 masks () generated by SAM [22] and 5k pixels (), with K = 16 to get smooth Gaussian features.
-6.2 Neu3D
+) on the objects of interest in the novel view to segment the target objects.9.2 Neu3D
 Each video contains 300 frames, and we begin by down-scaling the images by a factor of 2, resulting in a resolution of .
@@
 | Method | mIoU | mAcc | time | mIoU | mAcc | time | mIoU | mAcc | time | mIoU | mAcc | time | mIoU | mAcc | time |
-Ours (2%) | 0.8719 | 0.9826 | 0.4068 | 0.8368 | 0.9790 | 2.3117 | 0.8738 | 0.9923 | 2.2506 | 0.9154 | 0.9942 | 9.6083 | 0.9258 | 0.9904 | 1.6118 |
-SADG (10%) | 0.8707 | 0.9828 | 1.5782 | 0.8384 | 0.9794 | 11.8430 | 0.8792 | 0.9930 | 6.1878 | 0.9160 | 0.9939 | 71.2429 | 0.9283 | 0.9908 | 6.8211 |
-SADG (25%) | 0.8762 | 0.9832 | 3.9425 | 0.8461 | 0.9809 | 43.2682 | 0.8886 | 0.9935 | 18.8976 | 0.9180 | 0.9943 | 221.7642 | 0.9297 | 0.9907 | 19.0767 |
+| *TRASE (2%) | 0.8760 | 0.9829 | 0.4068 | 0.8368 | 0.9790 | 2.3117 | 0.8738 | 0.9923 | 2.2506 | 0.9154 | 0.9942 | 9.6083 | 0.9258 | 0.9904 | 1.6118 |
+| TRASE (10%) | 0.8707 | 0.9828 | 1.5782 | 0.8384 | 0.9794 | 11.8430 | 0.8792 | 0.9930 | 6.1878 | 0.9160 | 0.9939 | 71.2429 | 0.9283 | 0.9908 | 6.8211 |
+| TRASE (25%) | 0.8762 | 0.9832 | 3.9425 | 0.8461 | 0.9809 | 43.2682 | 0.8886 | 0.9935 | 18.8976 | 0.9180 | 0.9943 | 221.7642 | 0.9297 | 0.9907 | 19.0767 |
 Similar to Neu3D, the center camera (camera_0001) is used for testing, while the other cameras are used for training.
 We generate the initial point cloud using SfM across all the first frames, excluding the test camera, and apply the same hyper-parameters as in Neu3D in Sec.
-6.2.
+9.2.
 9.4 Technicolor Light Field
@@
 We generate the initial point cloud using SfM across all the first frames excluding the test camera and apply the same hyper-parameters as in Neu3D in Sec.
-6.2.
+9.2.
 10 Testing Implementation Details
-After the dynamic geometry reconstruction and the Gaussian features are properly trained, we proceed to our GUI to select the object of interest and evaluate the performance of the model by mIoU and mAcc as discussed in Sec.
-4.4.
+After the dynamic geometry reconstruction and Gaussian features are properly trained, we proceed to our GUI to select the object of interest and evaluate the model’s performance.
 Given a clicked point on the novel view, we first use the known camera intrinsics to compute the inverse intrinsic matrix .
@@
 3.
+Optional: Adjust the threshold to refine the object’s boundaries (, , , , ).
+4.
 Click the button ”Render Object” to render the selected object to test views for evaluating the performance.
-HyperNeRF-Mask | Neu3D-Mask | Immersive-Mask | Technicolor-Mask | |||||
-Method | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
-SA4D [16] | 0.7767 | 0.9633 | 0.8832* | 0.9391* | 0.7987* | 0.9835* | 0.8271* | 0.9734* |
-SADG (Ours) | 0.8663 | 0.9845 | 0.9022 | 0.9945 | 0.9234 | 0.9945 | 0.9308 | 0.9917 |
 11 Ablation Studies
-8.1 Filtering
-We noticed that the raw segmented object from the selected Gaussian clusters sometimes has some artifacts near the outline of the objects as illustrated in Fig.
-9.
-Therefore, we adopt a simple yet effective filtering step to filter out those artifacts.
-The procedure of the filtering is as follows:
-In each selected Gaussian cluster , we compute the mean Gaussian features in the selected cluster .
-We compute the per-Gaussian similarity score between the Gaussian features inside the selected cluster and mean Gaussian features of the selected cluster with cosine similarity.
-Delete Gaussians that have lower similarity scores than the threshold .
-As shown in Tab. 2, the proposed filtering step can effectively improve the average mIoU and mAcc for each benchmark.
 11.1 Smooth Gaussian Features
-In Sec.
-3.2 we introduced smooth Gaussian features (Eq.
-6).
+As mentioned in Sec. 3.2, we utilize smooth versions of Gaussian features, which can be defined as follows:
 We provide a detailed analysis of the selection of the number of nearest neighbors for the KNN algorithm in Tab.
-3.
+5.
 Overall, when smooth Gaussian features are enabled ( and ), the performance is better than when smooth Gaussian features are disabled ().
@@
 11.2 Gaussian Feature Clustering
-In Sec.
-3.2 we additionally introduced Gaussian Feature Clustering and the sub-sampling strategy to run the clustering in real-time with scenes that normally contain more than 300k Gaussians.
+We also employ a subsampling strategy to enable real-time Gaussian feature clustering in scenes that typically contain more than 300k Gaussians.
 We perform time analysis with different sub-sampling numbers of Gaussians (2%, 10%, and 25%) on RTX3080 and Intel i7-12700.
@@
 We report the average clustering time across all scenes in each benchmark and the corresponding mIoU and mAcc for our object of interest in Tab.
-4.
+6.
 Overall, with different sub-sampling of Gaussians (2%, 10%, and 25%), the mIoU and mAcc stay almost unchanged.
-However, as increasing the number Gaussians results in a longer clustering time, we believe that 2% is the optimal sub-sampling number for clustering.
+However, as increasing the number of Gaussians results in a longer clustering time, we believe that 2% is the optimal sub-sampling number for clustering.
+11.3 Sensitivity of Post-processing Filtering Threshold
+We also evaluated the sensitivity of the post-processing filtering threshold.
+As illustrated in Fig.
+13, there exists a wide margin of effective values.
+This demonstrates that selecting a threshold is not difficult, as one can easily choose a valid setting by observing the performance tendency.
 11.4 DBSCAN vs. K-Means
-There are various clustering techniques and K-means [29, 32, 15, 39] is one of the most widely used.
+There are various clustering techniques and K-means [27, 30] is one of the most widely used.
 K-means partitions data into clusters, where each feature belongs to the cluster with the nearest centroid (mean of the cluster).
@@
 As shown in Fig.
-10, the choice of significantly impacts the result of rendered segmentation.
+14, the choice of significantly impacts the result of rendered segmentation.
 Since different sequences contain varying numbers of objects, we believe that employing a method with an adaptive number of clusters, such as DBSCAN, validates our design choices with additional flexibility and no loss on the segmentation accuracy.
-8.5 Time and Storage Analysis
-We performed a comprehensive time analysis of SADG using an RTX 3080 GPU and an Intel i7-12700 CPU on the NeRF-DS [52] dataset.
-To evaluate its performance, we compared it with DGD [25], the only publicly available model addressing segmentation in dynamic novel views.
-The results are summarized in Tab.
-6.
-Most existing approaches for novel view segmentation [57, 6, 16] require precomputing SAM masks or object masks prior to training.
-In contrast, DGD dynamically computes DINOv2 feature maps for supervision during training.
-For a fair comparison, we precomputed the DINOv2 feature maps for all training images in DGD and stored them beforehand on an SSD, alongside our precomputed SAM masks.
+| NeRF-DS-Mask | HyperNeRF-Mask | Neu3D-Mask | ||||
+| Sampling rule | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
+| By Area | 0.8403 | 0.9743 | 0.8171 | 0.9632 | 0.8679 | 0.9920 |
+| By Score | 0.8028 | 0.9643 | 0.7998 | 0.9722 | 0.7673 | 0.9831 |
+| Grid-based | 0.8721 | 0.9822 | 0.8484 | 0.9809 | 0.8708 | 0.9922 |
+| Random (Ours) | 0.8768 | 0.9831 | 0.8663 | 0.9845 | 0.9022 | 0.9945 |
+11.5 Gaussian Feature Dimension
+We also analyze the impact of Gaussian feature dimensionality on segmentation performance.
 As shown in Tab.
-6, DGD [25] spends little preprocessing time for DINOv2 features (approximately 5 seconds for all training images per sequence).
-Although our SADG requires preprocessing time to generate SAM masks, this process only needs to be done once before training.
-DGD assigns a 384-dimensional semantic feature to each Gaussian.
-Rendering these high-dimensional features into 2D is computationally inefficient, significantly slowing down the training process.
-SADG uses 32-dimensional Gaussian features, which can be rendered much more efficiently.
-As a result, SADG achieves over 3 faster total time (preprocessing + training) than DGD while only requiring 4.5 less space in the storage.
-In addition, we evaluated the processing time for click prompts in our GUI.
-For smaller scenes with less than 200k Gaussians such as as_novel_view, segmenting an object via a click prompt takes approximately 0.05 seconds.
-For larger scenes with more than 1,000k Gaussians captured by a multi-view camera setup, such as 01_Welder, the processing time per click prompt is around 0.4 seconds.
-Avg. Preprocessing Time (min) | Avg. Training Time (min) | Total Time (min) | Avg. Size of Gaussians (MB) | |
-DGD [25] | 0.08 | 103 | 103 | 285 |
-SADG (Ours) | 16 | 16 | 32 | 60 |
-9 Comparison Against Segment Any 4D Gaussians Model
-We provide additional quantitative and qualitative comparisons with the concurrent work on 4D segmentation and editing, SA4D [16].
-The pipeline consists of three stages: (1) dynamic 3D reconstruction based on 4D-GS model [50], (2) identity feature field learning, and (3) post-processing with outlier removal.
-Since the authors have not released the official code, we use the unclean version from https://github.com/guanjunwu/sa4d for training and testing the model on our benchmark.
-We do not provide the performance on NeRf-DS-Mask with specular and transparent moving objects in Tab.
-5 since 4D-GS fails to converge on most of the sequences apart from two (basin_novel_view, press_novel_view).
-Similarly, we exclude two sequences from the HyperNeRF-Mask dataset, such as hand1-dense-v2 and espresso due to 4D reconstruction failure.
-In quantitative assessment, we select the top 30 segmentation results to more closely reflect the test set used in the original SA4D benchmark (not released).
-We also uniformly sub-sample point cloud (100K points) from SfM [45].
-As seen in Tab.
-5, our SADG significantly outperforms SA4D on all datasets.
-From the qualitative results in Fig.
-12 and Fig.
-13, we notice that rendered masks are larger than the underlying object and include a lot of unrelated regions even after the outlier removal in the post-processing step, which negatively affects objects’ segmentation accuracy.
-We further demonstrate some editing comparisons in Fig.
-17, Fig.
-16, and Fig.
-15.
-Due to inaccurate segmentation, composition, and removal, SA4D results contain many spurious regions, degrading the editing quality.
-10 Quantitative Results per Sequence
-For completeness to the average numbers per datasets (Tab. 1), we report the performance of our method individually on all benchmarks in Tab. 7, Tab. 8, Tab. 9, Tab. 10, and Tab. 11.
-11 Segmentation Benchmarks
-12 Additional Qualitative Results
-We provide additional qualitative results in Fig.
-12 and Fig.
-13 to further showcase the effectiveness of SADG over other methods.
-While SAGA [7] suffers from degenerate Gaussians, our reconstruction backbone and effective post-processing mitigate the issue.
-At the same time, rendered masks from other methods, such as DGD [25] and SA4D [16], are inaccurate and introduce spurious regions.
-Our method demonstrates the best performance among all approaches.
-13 Further Scene Editing Examples
-We present further scene composite examples of SADG in Fig.
-14.
-We also compare the quality of object removal and scene composition with SA4D [16] in Fig.
-16 and Fig.
-17.
-Overall, as discussed in Sec.
-9, SADG outperforms SA4D both quantitatively and qualitatively.
+8, increasing does not lead to a clear improvement in mIoU or mAcc.
+On the other hand, the runtime costs such as storage requirements and clustering time increase significantly with larger as shown in Tab.
+9.
+Fig.
+15 visualizes these trade-offs by plotting the no-filtered mIoU from HyperNeRF-Mask against runtime costs.
+Ideally, the optimal configuration should be located at the top-left corner (high mIoU, low cost).
+The plot confirms that our chosen dimensionality () strikes the best balance between segmentation performance and computational efficiency.
+| NeRF-DS-Mask | HyperNeRF-Mask | Neu3D-Mask | |||||
+| D | Filtering | mIoU | mAcc | mIoU | mAcc | mIoU | mAcc |
+| D=16 | ✘ | 0.8730 | 0.9822 | 0.8339 | 0.9778 | 0.8702 | 0.9921 |
+| ✓ | 0.8733 | 0.9822 | 0.8563 | 0.9828 | 0.9036 | 0.9945 | |
+| D=32 | ✘ | 0.8760 | 0.9829 | 0.8368 | 0.9790 | 0.8738 | 0.9923 |
+| ✓ (Ours) | 0.8768 | 0.9831 | 0.8663 | 0.9845 | 0.9022 | 0.9945 | |
+| D=64 | ✘ | 0.8744 | 0.9825 | 0.8347 | 0.9780 | 0.8949 | 0.9939 |
+| ✓ | 0.8756 | 0.9830 | 0.8573 | 0.9832 | 0.9068 | 0.9948 | |
+| D=128 | ✘ | 0.8752 | 0.9826 | 0.8320 | 0.9775 | 0.8936 | 0.9938 |
+| ✓ | 0.8761 | 0.9830 | 0.8573 | 0.9830 | 0.9070 | 0.9948 | |
+| NeRF-DS-Mask | HyperNeRF-Mask | Neu3D-Mask | ||||
+| D | Size | Time | Size | Time | Size | Time |
+| D=16 | 55 | 0.17 | 168 | 0.59 | 198 | 0.69 |
+| D=32 (Ours) | 67 | 0.30 | 203 | 1.10 | 239 | 1.56 |
+| D=64 | 90 | 0.43 | 272 | 4.12 | 320 | 3.46 |
+| D=128 | 135 | 0.82 | 410 | 8.90 | 483 | 7.06 |
+11.6 Mask Sampling
+We further study the influence of SAM [22] masks sampling during the feature learning stage.
+We test three different sampling heuristics: spatially uniform, based on the mask area (number of pixels), and based on the mask score.
+For the latter, we consider the stability scores from SAM [22] model, which reflects the reliability of the object against different object’s boundary changes.
+We have noticed many partial segments receiving high scores, especially for objects with reflective surfaces from the NeRF-DS dataset.
+This negatively affects the segmentation accuracy for the model that prioritizes the masks with the highest stability scores.
+Spatially uniform masks are retrieved by dividing the image space into 50 cells and sampling a mask per cell.
+If several candidates correspond to one cell, masks are selected based on their area.
+This method generally works well.
+However, true uniform sampling highly depends on mask shape, size, and arrangement, which is hard to achieve with simple heuristics.
+12 Limitations and Future Work
+We observe that TRASE faces difficulties in separating small objects that significantly overlap with larger ones, as illustrated in Fig.
+16.
+Intuitively, the scale-aware training proposed by SAGA [6] should improve the segmentation of these smaller objects; however, our experiments indicate that their pipeline yields noisy segmentation results.
+A potential alternative is to refine the construction of the mask-based correspondence matrix .
+While the current implementation encodes it with binary values (0 or 1), employing soft or probabilistic values could offer greater robustness.
+As described in the main paper, our method follows a two-stage training scheme.
+Thus, exploring the benefits of learning dynamic reconstruction and segmentation jointly is an exciting future direction of TRASE.
+Moreover, due to the rapid advances in the field, novel dynamic 3DGS reconstruction backbones can be combined with our soft-mining feature learning to achieve better segmentation performance.
+13 Additional Evaluations of Soft-Mining Contrastive Objective
+Soft vs.
+Hard vs.
+All Mining.
+To validate the effectiveness of our soft-mined pixel-pair sampling strategy, we compare its average intra-cluster and inter-cluster distances against two baselines: sampling all positive–negative pairs (all) and sampling only pairs that individually satisfy the thresholds (hard), as shown in Tab.
+10.
+Our results show that soft sample mining reduces intra-cluster distances while increasing inter-cluster distances, thereby improving feature separability.
+| NeRF-DS-Mask | HyperNeRF-Mask | Neu3D-Mask | ||||
+| Method | intra () | inter () | intra () | inter () | intra () | inter () |
+| TRASE (all) | 0.2588 | 1.3756 | 0.3251 | 1.3382 | 0.2200 | 1.3894 |
+| TRASE (hard) | 0.3517 | 1.1027 | 0.3488 | 1.1078 | 0.2942 | 1.2600 |
+| TRASE (Ours) | 0.2538 | 1.3761 | 0.2437 | 1.3406 | 0.1752 | 1.3925 |
+Comparison to Contrastive Methods.
+Existing contrastive segmentation methods CGC [12] and OpenGaussian [46] treat all negative pairs equally and focus on average feature consistency.
+However, this often leads to ambiguous feature assignments (Fig.
+17) and segmentation artifacts (Fig.
+18).
+In contrast, TRASE optimizes per-pixel relationships by introducing soft sample mining as discussed in our main paper.
+It prioritizes challenging pixels that have any negative pixel pairs that are close in feature space and any positive pairs that are far apart, leading to more discriminative feature representations.
+We confirm our findings by evaluating segmentation accuracy in Tab.
+1 (main).
+Our TRASE achieves the best segmentation performance across all tested benchmarks, validating the effectiveness of the learned segmentation field.
+14 Additional Feature Space Visualizations
+As shown in Fig.
+17, TRASE demonstrates superior performance in separating different clusters apart compared to the existing contrastive learning approaches [46, 42].
+We provide additional results in Fig.
+20.
+The features learned by TRASE are more discriminative compared to other approaches.
+15 Additional Out-of-FoV Visualizations
+As shown in the main paper, TRASE with soft sample mining demonstrates superior performance in the out-of-FoV viewpoint to eliminate floating artifacts compared to TRASE with hard / all sample mining and the existing contrastive learning approaches [46, 42].
+We provide additional results in Fig.
+18.
 | as_novel_view | basin_novel_view | cup_novel_view | press_novel_view | plate_novel_view | Average | |||||||
@@
 | SAGA [6] | 0.8235 | 0.9542 | 0.6437 | 0.9140 | 0.8833 | 0.9922 | 0.8119 | 0.9918 | 0.8736 | 0.9699 | 0.8072 | 0.9644 |
+| SA4D [19] | 0.7384 | 0.9158 | 0.7663 | 0.9609 | 0.7759 | 0.9858 | 0.6558 | 0.9894 | 0.4337 | 0.8282 | 0.6740 | 0.9360 |
 | DGD [24] | 0.7135 | 0.9266 | 0.6685 | 0.9291 | 0.694 | 0.9752 | 0.6457 | 0.9822 | 0.8406 | 0.9625 | 0.7125 | 0.9551 |
-SADG (Ours) | 0.9162 | 0.9799 | 0.8740 | 0.9785 | 0.8864 | 0.9924 | 0.8098 | 0.9916 | 0.8733 | 0.9707 | 0.8719 | 0.9826 |
+| CGC [12] | 0.7388 | 0.9404 | 0.8084 | 0.9659 | 0.8556 | 0.9908 | 0.7723 | 0.9905 | 0.8748 | 0.9716 | 0.8100 | 0.9718 |
+| OpenGaussian [46] | 0.7471 | 0.9438 | 0.8580 | 0.9759 | 0.8406 | 0.9896 | 0.3047 | 0.9500 | 0.7191 | 0.9228 | 0.6939 | 0.9564 |
+| TRASE (Ours) | 0.9160 | 0.9798 | 0.8704 | 0.9775 | 0.8893 | 0.9927 | 0.8257 | 0.9926 | 0.8826 | 0.9729 | 0.8768 | 0.9831 |
 | americano | chickchicken | cut-lemon1 | espresso | hand | keyboard | |||||||
@@
 | SAGA [6] | 0.7325 | 0.9873 | 0.8633 | 0.9648 | 0.5263 | 0.8398 | 0.6395 | 0.9798 | 0.9088 | 0.9859 | 0.8070 | 0.9645 |
+| SA4D [19] | 0.8488 | 0.9941 | 0.8269 | 0.9553 | 0.8161 | 0.9688 | 0.4507 | 0.9640 | 0.7064 | 0.9462 | 0.7930 | 0.9651 |
 | DGD [24] | 0.7748 | 0.9912 | 0.8317 | 0.9589 | 0.8583 | 0.9718 | 0.7240 | 0.9871 | 0.8936 | 0.9838 | 0.8451 | 0.9733 |
-SADG (Ours) | 0.8144 | 0.9922 | 0.9308 | 0.9835 | 0.8795 | 0.9769 | 0.7164 | 0.986 | 0.9006 | 0.9849 | 0.8952 | 0.9827 |
+| CGC [12] | 0.7917 | 0.9915 | 0.8625 | 0.9672 | 0.8047 | 0.9605 | 0.7586 | 0.9898 | 0.7692 | 0.9646 | 0.8838 | 0.9810 |
+| OpenGaussian [46] | 0.5739 | 0.9723 | 0.8309 | 0.9583 | 0.8369 | 0.9659 | 0.5949 | 0.9760 | 0.7687 | 0.9618 | 0.3848 | 0.8051 |
+| TRASE (Ours) | 0.8144 | 0.9922 | 0.9308 | 0.9835 | 0.8795 | 0.9769 | 0.7164 | 0.986 | 0.9006 | 0.9849 | 0.8952 | 0.9827 |
 | oven-mitts | slice-banana | split-cookie | torchocolate | Average | ||||||||
@@
 | SAGA [6] | 0.8667 | 0.9766 | 0.7323 | 0.8960 | 0.8186 | 0.9906 | 0.7652 | 0.9933 | 0.7660 | 0.9579 | ||
-DGD [25] | 0.8595 | 0.9755 | 0.8408 | 0.947 | 0.8467 | 0.9925 | 0.8223 | 0.9954 | 0.8297 | 0.9777 | ||
-SADG (Ours) | 0.9295 | 0.9876 | 0.8782 | 0.9617 | 0.8581 | 0.9930 | 0.8599 | 0.9964 | 0.8663 | 0.9845 |
+| SA4D [19] | 0.7257 | 0.9404 | 0.7205 | 0.8983 | 0.8351 | 0.9931 | 0.6474 | 0.9911 | 0.7371 | 0.9616 | ||
+| DGD [24] | 0.8595 | 0.9755 | 0.8408 | 0.9470 | 0.8467 | 0.9925 | 0.8223 | 0.9954 | 0.8297 | 0.9777 | ||
+| CGC [12] | 0.9025 | 0.9832 | 0.8378 | 0.9469 | 0.7975 | 0.9906 | 0.7488 | 0.9929 | 0.8157 | 0.9768 | ||
+| OpenGaussian [46] | 0.9107 | 0.9845 | 0.7470 | 0.9170 | 0.8816 | 0.9946 | 0.0011 | 0.9440 | 0.6531 | 0.9480 | ||
+| TRASE (Ours) | 0.9295 | 0.9876 | 0.8782 | 0.9617 | 0.8581 | 0.9930 | 0.8599 | 0.9964 | 0.8663 | 0.9845 | ||
 | coffee_martini | cook_spinach | cut_roasted_beef | flame_steak | sear_steak | Average | |||||||
@@
 | SAGA [6] | 0.2201 | 0.8081 | 0.8125 | 0.9881 | 0.6982 | 0.9767 | 0.8439 | 0.9911 | 0.8959 | 0.9941 | 0.6941 | 0.9516 |
-DGD [25] | 0.7875 | 0.9865 | 0.815 | 0.9883 | 0.817 | 0.9877 | 0.6771 | 0.9776 | 0.7638 | 0.9854 | 0.7721 | 0.9851 |
-SADG (Ours) | 0.9120 | 0.9948 | 0.9129 | 0.9951 | 0.9103 | 0.9947 | 0.8716 | 0.9930 | 0.9044 | 0.9947 | 0.9022 | 0.9945 |
+| SA4D [19] | 0.8583 | 0.9910 | 0.8987 | 0.9941 | 0.8645 | 0.9914 | 0.8898 | 0.9940 | 0.9047 | 0.9948 | 0.8832 | 0.9931 |
+| DGD [24] | 0.7875 | 0.9865 | 0.8150 | 0.9883 | 0.8170 | 0.9877 | 0.6771 | 0.9776 | 0.7638 | 0.9854 | 0.7721 | 0.9851 |
+| CGC [12] | 0.8359 | 0.9895 | 0.9034 | 0.9945 | 0.9052 | 0.9943 | 0.8530 | 0.9920 | 0.8794 | 0.9934 | 0.8754 | 0.9927 |
+| OpenGaussian [46] | 0.8254 | 0.9896 | 0.6336 | 0.9798 | 0.9115 | 0.9951 | 0.8199 | 0.9907 | 0.8986 | 0.9943 | 0.8178 | 0.9899 |
+| TRASE (Ours) | 0.9120 | 0.9948 | 0.9129 | 0.9951 | 0.9103 | 0.9947 | 0.8716 | 0.9930 | 0.9044 | 0.9947 | 0.9022 | 0.9945 |
 | 01_Welder | 02_Flames | 10_Alexa_1 | 11_Alexa_2 | Average | ||||||
@@
 | SAGA [6] | 0.8062 | 0.9737 | 0.8463 | 0.9940 | 0.4794 | 0.9513 | 0.8260 | 0.9797 | 0.7395 | 0.9747 |
+| SA4D [19] | 0.7728 | 0.9708 | 0.7200 | 0.9871 | 0.8351 | 0.9907 | 0.8669 | 0.9853 | 0.7987 | 0.9835 |
 | DGD [24] | 0.873 | 0.9854 | 0.7775 | 0.9914 | 0.8107 | 0.9899 | 0.7312 | 0.9734 | 0.7981 | 0.9850 |
-SADG (Ours) | 0.8975 | 0.9877 | 0.8808 | 0.9955 | 0.9367 | 0.9970 | 0.9786 | 0.9979 | 0.9234 | 0.9945 |
+| CGC [12] | 0.911 | 0.9894 | 0.8520 | 0.9947 | 0.8464 | 0.9924 | 0.9328 | 0.9933 | 0.8856 | 0.9925 |
+| OpenGaussian [46] | 0.9077 | 0.9891 | 0.1439 | 0.9419 | 0.8752 | 0.9941 | 0.9213 | 0.9922 | 0.7120 | 0.9793 |
+| TRASE (Ours) | 0.8975 | 0.9877 | 0.8808 | 0.9955 | 0.9367 | 0.9970 | 0.9786 | 0.9979 | 0.9234 | 0.9945 |
 | Birthday | Fabien | Painter | Theater | Average | ||||||
@@
 | SAGA [6] | 0.8541 | 0.9779 | 0.9371 | 0.9943 | 0.9670 | 0.9867 | 0.4071 | 0.9579 | 0.7913 | 0.9792 |
+| SA4D [19] | 0.8322 | 0.9736 | 0.8709 | 0.9460 | 0.8070 | 0.9808 | 0.7984 | 0.9930 | 0.8271 | 0.9734 |
 | DGD [24] | 0.7203 | 0.9573 | 0.9177 | 0.9667 | 0.7959 | 0.9788 | 0.6826 | 0.9885 | 0.7791 | 0.9728 |
-SADG (Ours) | 0.8839 | 0.9833 | 0.9734 | 0.9889 | 0.9719 | 0.9976 | 0.8941 | 0.9969 | 0.9308 | 0.9917 |
+| CGC [12] | 0.7951 | 0.9668 | 0.9163 | 0.9659 | 0.9668 | 0.9971 | 0.8443 | 0.9950 | 0.8806 | 0.9812 |
+| OpenGaussian [46] | 0.7008 | 0.9458 | 0.8590 | 0.9438 | 0.8928 | 0.9908 | 0.9034 | 0.9971 | 0.8390 | 0.9694 |
+| TRASE (Ours) | 0.8839 | 0.9833 | 0.9734 | 0.9889 | 0.9719 | 0.9976 | 0.8941 | 0.9969 | 0.9308 | 0.9917 |
+16 Further Scene Editing Examples
+We present further scene composite examples of TRASE in Fig.
+24.
+We also compare the quality of object removal and scene composition with SA4D [19] in Fig.
+25 and Fig.
+26.
+While SA4D introduces ghostly artifacts in place of the removed objects and carries spurious Gaussians with the segmented objects, our TRASE delivers photorealistic results of the edited scene.
+17 Detailed Quantitative Results per Dataset
+18 Additional Qualitative Results
+We provide additional qualitative results in Fig.
+21 and Fig.
+22 to further showcase the effectiveness of TRASE over other methods.
+While SAGA [7] suffers from degenerate Gaussians, our reconstruction backbone and effective post-processing mitigate the issue.
+At the same time, rendered masks from other methods, such as DGD [24] and SA4D [19], are inaccurate and introduce spurious regions.
+Our method demonstrates the best performance among all approaches.
+We also showcase in Fig.
+23 that TRASE achieves both time and space consistency in multiview video dataset such as Neu3D.
 19 Visualization of the Rendered View, Rendered Segmentation, Segmented Object
 We also provide the visualizations of the rendered novel view, the rendered segmentation, and the segmented objects in Fig.
-18, Fig.
-19, and Fig.
-20.
-The rendered segmentation of SADG is based on Gaussian clustering, which is performed in 3D space, resulting in consistent segmentation in the dynamic scene.
+27, Fig.
+28, and Fig.
+29.
+The rendered segmentation of TRASE is based on Gaussian clustering, which is performed in 3D space, resulting in consistent segmentation in the dynamic scene.